### v1

# ConvTimeNet: A Deep Hierarchical Fully Convolutional Model for Multivariate Time Series Analysis

Abstract
This paper introduces ConvTimeNet, a novel deep hierarchical fully convolutional network designed to serve as a general-purpose model for time series analysis. The key design of this network is twofold, designed to overcome the limitations of traditional convolutional networks. Firstly, we propose an adaptive segmentation of time series into sub-series level patches, treating these as fundamental modeling units. This setting avoids the sparsity semantics associated with raw point-level time steps. Secondly, we design a fully convolutional block by skillfully integrating deepwise and pointwise convolution operations, following the advanced building block style employed in Transformer encoders. This backbone network allows for the effective capture of both global sequence and cross-variable dependence, as it not only incorporates the advancements of Transformer architecture but also inherits the inherent properties of convolution. Furthermore, multi-scale representations of given time series instances can be learned by controlling the kernel size flexibly. Extensive experiments are conducted on both time series forecasting and classification tasks. The results consistently outperformed strong baselines in most situations in terms of effectiveness. The code is publicly available. [^1] https://github.com/Mingyue-Cheng/ConvTimeNet.
1 Introduction
Time series is a sequence of data points arranged in chronological order. Fully analyzing and mining this type of data Roberts et al. (2013); Kim et al. (2021), such as forecasting evolving trends or classifying the target labels for a given instance, can yield significant gains in various fields, including economics, finance, and biology. Over the past decades, numerous efforts Box et al. (2015); Liu et al. (2021b) have been dedicated to this area. Among these efforts, deep learning-based methods Zhang et al. (2020); Foumani et al. (2023); Liu et al. (2024a) have made significant progress due to their ability to model non-linear relationships and eliminate labor-intensive feature engineering.
Over a significant period in the past, the convolutional network He et al. (2016); Zheng et al. (2014); Middlehurst et al. (2023) has played a crucial role in time series analysis, largely due to its inherent properties that strike an excellent balance between computational efficiency and representation quality. Data back to the past years, many representative works Bagnall et al. (2017) of time series analysis typically employ convolutional networks as the backbone. For instance, temporal convolutional network (TCNBai et al. (2018)) and its variants are widely used in modeling temporal variation dependence for the time series forecasting task. Furthermore, a large number of works (such as InceptionTimeIsmail Fawaz et al. (2020), MiniRocketDempster et al. (2021), and MCNNCui et al. (2016)) are also proposed by employing convolutional networks to identify informative patterns from given instances in the classification of time series. However, with the advent of Transformer networks Vaswani et al. (2017); Dosovitskiy et al. (2020), the role of convolutional architecture in time series analysis appears to be diminishing. Recently, Transformer-style networks Wen et al. (2022) have nearly become the dominant role in learning representations of time series, ranging from forecasting Zhou et al. (2021); Liu et al. (2023) to classification tasks Liu et al. (2021a).
In reality, the primary motivation for widely employing the Transformer network is its utilization of the self-attention mechanism. This mechanism enables a more effective capture of long-term sequence dependencies. Despite their effectiveness, using Transformer networks for time series analysis also has some areas of incompatibility. Firstly, multi-scale representations, a crucial element in time series analysis, are challenging to learn within the vanilla Transformer network. This is because the self-attention mechanism can only model fixed-scale representations of sequence input. Secondly, unlike other types of sequence data, both temporal variation Xiao et al. (2022) and cross-variable dependence modeling Fang et al. (2023) are crucial for time series analysis. However, the Transformer network lacks sufficient consideration of modeling interactions across variables, which carry crucial information for accurate predictions of time series. Thirdly, it usually produces quadratic computation complexity Zhou et al. (2021) while attending very long point-wise sequence inputs. This is difficult to avoid because a numerical value at each time step can struggle to reveal informative information, resulting in very long sequence input.
Through careful analysis, we observe that these aforementioned predicaments can largely be avoided by leveraging the inherent properties of convolutional networks. Specifically, firstly, the ability to learn varying scale representations is easily achieved by maintaining a hierarchical convolutional network architecture Cui et al. (2016). Secondly, it is established that the convolutional filter operator naturally complements the modeling of cross-variable relationships Liu et al. (2023). When it comes to temporal dependence, the receptive field can be progressively expanded by stacking deep convolutional operations, thereby acquiring the capacity for global sequence dependence. Thirdly, convolutional neural networks are highly efficient in computation, primarily due to their weight-sharing mechanism Han et al. (2021). This efficiency distinguishes them from Transformer networks, which are constrained by the curse of quadratic time complexity, which is directly proportional to the square of the input size. Along this line, two questions naturally arise: what is the primary obstacle in applying current convolutional networks to time series analysis? If these barriers are overcome, can convolutional networks be revolutionized to achieve superior performance again?
In response to the aforementioned questions, in this work, we propose a novel convolutional network, dubbed ConvTimeNet, designed to serve as a versatile model for time series analysis. One of the most distinctive features of ConvTimeNet is its use of pure convolutional operators to learn the representation of given time series data. Crucially, the network not only preserves the advanced properties of Transformer networks but also inherits several inherent strengths of convolutional networks. Specifically, we first adhere to the modern philosophy of neighboring sequence points together to avoid the sparsity semantic dilemma carried by a single numerical value. Notably, we propose to divide the raw time series into sequences of patches in a data-driven manner, which can better preserve the semantics of local regions. Secondly, on the basis of the building block in Transformer encoder networks, we design a novel fully convolutional block, where deepwise and pointwise convolutions are organized together simultaneously. Particularly, the proposed ConvTimeNet highlights that very deep and hierarchical network architectures are encouraged and very significant in modeling global receptive fields and learning multi-scale representations of given time series instances. To demonstrate the effectiveness and generic capacity of ConvTimeNet, we conduct extensive experiments over a variety of time series analysis tasks including forecasting and classification. The experimental results show that ConvTimeNet could achieve superior or competitive performance compared to strong baselines, including both advanced Transformer networks and pioneering convolutional models. We hope that the proposed ConvTimeNet can serve as an alternative and useful encoder for time series analysis.
2 A Recipe of Fully Convolutional Network for Time Series Analysis
Figure 1: Illustration of the newly proposed ConvTimeNet.
In this section, we begin by providing a brief overview of the architecture of the newly proposed ConvTimeNet. Subsequently, we delve into two key characteristics: deformable patch embedding and fully convolutional blocks. Finally, we demonstrate the strengths of the proposed ConvTimeNet, examining its capabilities from three distinct perspectives.
2.1 Overall Network Architecture
As illustrated in Figure 1, the forward pass of ConvTimeNet begins with the input of multivariate time series data. Before learning the representation of input data with hidden layers, we develop a deformable patch embedding layer. This layer’s primary role is to transform the raw point-wise time series input into a sequence of patch embeddings. Specifically, the selection of time series points for each patch is determined in a data-driven manner, facilitated by a predictor layer. Subsequently, the transformed time series representation is fed into a stack of full-convolution blocks, with the objective of capturing long-term dependence and learning multi-scale representations. Finally, the architecture culminates in a linear layer, designed to perform time series mining tasks such as forecasting or classification of given time series data.
2.2 Deformable Patch Embedding
The primary objective of time series analysis is to discern the correlation between data points across different time steps. In most existing studies, preserving raw point-wise time series data as input is nearly a default practice. However, a single-time step does not carry as much semantic meaning as a word in a sentence. Consequently, enhancing the information density of each input unit is crucial for improving the final results. We also note that some recent Transformer studies have begun to attempt to extract local information while analyzing their correlations. A common strategy is to aggregate neighboring time points into sub-series level patches. This process, in contrast to modeling each point-wise input individually, actually groups time points into local regions, thereby coupling their dependencies. While this approach is effective, we argue that the fixed-size patch splitting may compromise semantics, leading to reduced performance. For instance, the same pattern in different time series might have varying sizes, a concept supported by well-known shapelet-based approaches Ye and Keogh (2009). Therefore, such strict patch splitting may capture inconsistent information across different instances, potentially leading to sub-optimal results.
Based on the preceding discussion, we assert that time series analysis has experienced success through patch splitting, yet substantial room for improvement remains. To address the dilemma mentioned above, we decided to adopt a philosophy of neighboring time points together, selecting specific time points within each patch in a data-driven manner. This idea of this approach is also consistent with previous work Dai et al. (2017); Chen et al. (2021). As illustrated in the lower left of Figure 1, we additionally introduced a lightweight predictor to adaptively select time points based on given vanilla features, followed by an embedding projection layer. In this study, we refer to this component as the DePatch module.
To achieve this goal, the DePatch module comprises three key steps. Formally, we denote the input time series as $X\in\mathbb{R}^{C\times T}$, where $C$ and $T$ indicate the number of variables (channels) and sequence lengths respectively. Before performing adaptive patch splitting, we first process the raw time series points into $N=\lfloor{\frac{T-P}{S}}\rfloor+2$ fixed-size patches, where the shape of each patch can be denoted by $x\in\mathbb{R}^{P\times C}$. We use the padding operation to ensure the uniform split. For clarity, we omit the superscript ($i$). Here, $P$ denotes the corresponding patch size while $S$ indicates the stride size of the patch window. Subsequently, each patch is assigned three key variables: the center position of this patch $x_{c}$, center offset $\delta_{c}$, and scale variation $\delta_{p}$. It should be noted that $x_{c}$ is decided by the uniform patch splitting above while the $\delta_{c}$ and $\delta_{p}$ are obtained by borrowing the following computation process:
$\delta_{c},\delta_{p}=H(g(x)),$ (1)
where $g(x)$ denotes the feature map obtained from input patch $x$, and $H(\cdot)$ indicates the lightweight predictor function. In practice, the lightweight predictor function can be instantiated by a convolutional-based projection layer. Particularly, extensive experiments over the experimental part show its sensitivity to the final results. Along this line, the final center of patch $x$ and its scale can be computed by
$x_{c}^{new}=x_{c}+\delta_{c}$ $,~{}~{}P^{new}=P+2*\delta_{p},$ (2)
$L=x_{c}^{new}-\frac{P^{new}}{2}$ $,~{}~{}R=x_{c}^{new}+\frac{P^{new}}{2},$
where $x_{c}^{new}$ denoted as the new center position of the patch, and $P^{new}$ represents the new length of patch. $L$ and $R$ are denoted as the left and right boundary positions. Finally, based on these two positions, we perform sampling on the original sequence with linear interpolation and obtain a new patch, followed by the projection layer transforms the new patch into the embedding space. With these efforts, the DePatch module allows for the adjustment of each patch’s position and scale based on the input features adaptively, thereby mitigating the semantic destruction caused by hard splitting.
2.3 Fully Convolutional Blocks
To learn high-quality data representation, past methods He et al. (2016); Vaswani et al. (2017) stacked repeated building blocks, a paradigm that is also very popular in the field of data representation learning. Typically, the Transformer block, as a classical design, has been successfully applied in various fields, and these successful applications have inspired methods in the field of time series. However, the Transformer block also has some drawbacks. Firstly, the traditional Transformer block mainly process fixed-scale input, which is not flexible enough. Secondly, the Transformer block is struggling to model the cross-variable dependencies. Finally, the computational complexity of the Transformer block is quadratic related to the length of the input sequence.
Luckily, convolution can effectively address these issues. Motivated by this, we propose a fully convolutional block, as shown in the lower right module of Figure 1. Specifically, we retain the Transformer block’s innovative architecture, incorporating linear and non-linear transformations. The primary distinction lies in our utilization of temporal dependency modeling, achieved through stacking 1D kernel groups, instead of the traditional multi-head attention mechanism. Practically, we employ deepwise convolution as a 1D kernel group, as it offers efficient computational efficiency. Meanwhile we substitute the Feed-Forward Network(FFN) in the Transformer block with a 1x1-sized pointwise convolution to achieve cross-variable dependency modeling. The synergistic combination of these two convolutions effectively completing the comprehensive modeling of the temporal and cross variable dependency.
To achieve a global receptive field, we increase the depth of the model. We discard the traditional residual connection and instead adopt a learnable residual Bachlechner et al. (2021), to circumvent the overfitting issue that often arises from an increased model depth. We also verifiy the effectiveness of the learnable residual mechanism in our experiments. Concurrently, we utilize Batch NormalizationIoffe and Szegedy (2015) in fully convolution block. Furthermore, we employ larger kernels to expand the receptive field. However, in practice, we found that training with large kernels presents a greater challenge. This observation is consistent with recent research findingDing et al. (2022). Therefore, we introduce a re-parameterization mechanism to address this issue. Due to the flexibility of convolutional networks, we can easily achieve a hierarchical architecture by altering the kernel size. We divide the entire deep network architecture into multiple distinct stages, each stage is composed of $K$ fully convolutional blocks. Drawing upon this hierarchical architecture, the time series representation across various scales can be effectively extracted. We also explored the impact of different configurations of hierarchical structures on the model’s capabilities in the experiment.
Concretely, denote $F_{\bm{\theta}}^{l}({\cdot})$ as the larger kernel branch and $G_{\bm{\theta}}^{l}({\cdot})$ as the small kernel branch in $l$-th fully convolution block. Note that $\theta$ represents the weight of the convolution kernel, it will be updated during gradient backpropagation. The re-parameterization mechanism is defined as follows:
${Z}_{DW}^{l}={Gelu}(F_{\bm{\theta}}^{l}({Z^{l-1}})+G_{\bm{\theta}}^{l}({Z^{l-1}})),$ (3)
where $Z^{l-1}$denote the output of $l-1$ convolutional block, $Z_{DW}^{l}\in\mathcal{\mathbb{R}}^{D\times M}$ represents the output deepwise convolution layer in the $l$-th fully convolutional block. $D$ denoted as the number of hidden dimensions and $M$ represents the new sequence length, which is equal to the number of patches $N$. During the inference phase in ConvTimeNet, a sophisticated technique is employed where the weights of the small kernel are zero padded and then combined with those of the larger kernel to form a new, unified convolutional kernel. Denote $\alpha$ as the learnable weights for the residual connections in deepwise convolution layer, typically initialized with 00.
$\hat{Z}_{DW}^{l}=Z^{l-1}+\alpha\times Z_{DW}^{l},$ (4)
where $\hat{Z}_{DW}^{l}$ denote the output of deepwise convolution in $l$-th building block. To sum up, by utilizing fully convolutional blocks, we can effectively model cross-dependencies within a single block. The design of hierarchical and large kernels for ConvTimeNet facilitates the representation learning of varying scales and the creation of a global receptive field.
2.4 Analysis of the ConvTimeNet
Through these efforts, a deep and hierarchical convolutional network has been constructed, serving as a general-purpose model for time series analysis. It is noteworthy that the proposed ConvTimeNet not only retains the advanced properties of the Transformer encoder but also inherits the pioneering convolutional network. Consequently, we explore the relationship between our ConvTimeNet and traditional convolutional networks, as well as Transformer networks.
ConvTimeNet is equipped with more modern techniques compared to traditional convolutional networks. It enhances the semantic information of the input through the use of a deformable patch module and incorporates numerous great designs from the Transformer architecture. Furthermore, unlike past convolutional methods in the time series field, we consider the convolution block from operator level and utilize deepwise convolution to enhance the computational efficiency of the model.
In comparison to the Transformer method, ConvTimeNet foregoes the complex self-attention module and instead constructs a cross-variable dependency model through the collaborative role of deepwise and pointwise convolutions. Simultaneously, due to the natural advantages of convolution, a multi-scale architecture can be directly implemented by varying the convolution kernels. The hierarchical architecture and large kernel in ConvTimeNet endow it with the capacity to possess a global receptive field.
Table 1: Experimental results of time series forecasting task evaluated by MSE and MAE.
Methods ConvTimeNet PatchTST iTransformer TiDE MSGNet TimesNet Crossformer DLinear NHits
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.368 0.394 0.385 0.408 0.405 0.419 0.374 0.395 0.423 0.440 0.421 0.438 0.390 0.417 0.375 0.396 0.423 0.444
192 0.406 0.414 0.419 0.426 0.448 0.447 0.409 0.417 0.465 0.469 0.482 0.479 0.424 0.448 0.428 0.437 0.504 0.493
336 0.405 0.420 0.429 0.434 0.482 0.470 0.435 0.433 0.468 0.473 0.528 0.505 0.486 0.492 0.448 0.449 0.513 0.503
720 0.442 0.457 0.446 0.462 0.560 0.537 0.446 0.460 0.540 0.524 0.527 0.510 0.507 0.519 0.505 0.514 0.622 0.564
ETTh2 96 0.264 0.330 0.278 0.341 0.305 0.361 0.290 0.350 0.348 0.399 0.355 0.408 0.803 0.628 0.296 0.360 0.318 0.373
192 0.316 0.368 0.343 0.382 0.391 0.412 0.349 0.388 0.404 0.431 0.403 0.434 1.028 0.743 0.391 0.423 0.425 0.447
336 0.315 0.378 0.372 0.404 0.418 0.433 0.371 0.409 0.375 0.419 0.398 0.434 1.167 0.828 0.445 0.460 0.596 0.527
720 0.382 0.425 0.395 0.430 0.437 0.455 0.401 0.437 0.421 0.451 0.443 0.465 1.665 1.032 0.700 0.592 1.353 0.810
ETTm1 96 0.292 0.345 0.298 0.345 0.306 0.360 0.310 0.352 0.309 0.362 0.331 0.372 0.345 0.394 0.303 0.346 0.323 0.376
192 0.329 0.368 0.339 0.374 0.345 0.382 0.345 0.372 0.356 0.392 0.435 0.421 0.461 0.483 0.338 0.368 0.365 0.401
336 0.363 0.390 0.381 0.401 0.378 0.402 0.379 0.391 0.393 0.414 0.457· 0.445 0.623 0.586 0.373 0.389 0.400 0.423
720 0.427 0.428 0.428 0.431 0.443 0.439 0.435 0.423 0.440 0.445 0.526 0.481 0.673 0.593 0.428 0.423 0.463 0.463
ETTm2 96 0.167 0.257 0.174 0.261 0.174 0.266 0.167 0.257 0.188 0.273 0.190 0.276 0.330 0.401 0.170 0.264 0.212 0.290
192 0.222 0.295 0.238 0.307 0.247 0.315 0.223 0.294 0.246 0.315 0.244 0.311 0.623 0.543 0.233 0.311 0.270 0.330
336 0.276 0.329 0.293 0.346 0.292 0.343 0.277 0.329 0.301 0.347 0.302 0.349 0.887 0.637 0.298 0.358 0.340 0.377
720 0.358 0.381 0.373 0.401 0.375 0.395 0.371 0.386 0.407 0.411 0.406 0.406 0.844 0.640 0.423 0.437 0.444 0.444
Electricity 96 0.132 0.226 0.138 0.233 0.132 0.228 0.160 0.262 0.169 0.279 0.177 0.281 0.150 0.258 0.141 0.238 0.147 0.250
192 0.148 0.241 0.153 0.247 0.154 0.249 0.174 0.275 0.188 0.296 0.193 0.295 0.175 0.284 0.154 0.251 0.154 0.261
336 0.165 0.259 0.170 0.263 0.172 0.267 0.190 0.289 0.199 0.307 0.206 0.306 0.218 0.325 0.170 0.269 0.177 0.280
720 0.205 0.293 0.206 0.295 0.204 0.296 0.229 0.319 0.227 0.330 0.223 0.320 0.226 0.324 0.205 0.302 0.211 0.313
Exchange 96 0.086 0.204 0.094 0.216 0.099 0.225 0.107 0.233 0.167 0.303 0.166 0.305 0.283 0.393 0.087 0.214 0.126 0.256
192 0.184 0.303 0.191 0.311 0.206 0.329 0.201 0.323 0.302 0.402 0.303 0.413 1.087 0.804 0.164 0.298 0.381 0.455
336 0.341 0.421 0.343 0.427 0.370 0.448 0.351 0.432 0.581 0.556 0.445 0.511 1.367 0.905 0.333 0.437 0.576 0.577
720 0.879 0.701 0.888 0.706 0.963 0.746 0.940 0.735 2.229 1.079 1.389 0.899 1.546 0.987 0.988 0.749 2.831 1.192
Traffic 96 0.376 0.265 0.395 0.272 0.361 0.266 0.445 0.325 0.567 0.337 0.600 0.323 0.514 0.292 0.411 0.284 0.436 0.313
192 0.392 0.271 0.411 0.278 0.378 0.271 0.458 0.331 0.579 0.339 0.612 0.327 0.528 0.298 0.423 0.289 0.449 0.317
336 0.405 0.277 0.424 0.284 0.390 0.274 0.471 0.336 0.604 0.350 0.628 0.344 0.538 0.303 0.437 0.297 0.475 0.330
720 0.436 0.294 0.453 0.300 0.424 0.291 0.500 0.352 0.637 0.359 0.657 0.349 0.559 0.312 0.467 0.316 0.517 0.350
Weather 96 0.155 0.205 0.147 0.197 0.162 0.212 0.178 0.229 0.160 0.217 0.168 0.225 0.148 0.214 0.176 0.236 0.156 0.212
192 0.200 0.249 0.191 0.240 0.205 0.251 0.221 0.264 0.208 0.255 0.218 0.268 0.201 0.270 0.217 0.275 0.203 0.260
336 0.252 0.287 0.244 0.282 0.257 0.291 0.268 0.298 0.272 0.302 0.269 0.301 0.248 0.311 0.264 0.315 0.266 0.319
720 0.321 0.335 0.320 0.334 0.325 0.337 0.336 0.345 0.357 0.356 0.340 0.350 0.366 0.395 0.325 0.364 0.338 0.365
Illness 24 1.469 0.800 1.657 0.869 1.930 0.883 3.988 1.444 2.453 1.044 2.130 0.981 4.312 1.418 2.313 1.059 2.575 1.146
36 1.450 0.845 1.467 0.813 1.807 0.904 3.937 1.432 2.628 1.021 2.312 1.013 4.034 1.319 2.402 1.102 2.676 1.128
48 1.572 0.875 1.833 0.913 1.894 0.945 4.074 1.451 2.808 1.106 2.334 1.042 4.303 1.393 2.420 1.111 2.607 1.135
60 1.741 0.920 2.168 1.009 2.033 1.013 3.981 1.434 3.008 1.157 2.051 0.989 0.390 0.417 2.400 1.116 2.531 1.087
Table 2: Experimental results of time series classification task evaluated by Accuracy, where full results in terms of both metrics of Accuracy and F1 score is shown in the supplement materials. Noted that -- indicates that the model cannot run due to the issue of being out of memory.
Datasets ConvTimeNet FormerTime TimesNet MiniRocket TST MLP TCN MCNN MCDCNN
AWR 0.987 0.978 0.980 0.972 0.980 0.960 0.884 0.977 0.980
CR 0.986 0.917 0.889 0.981 0.898 0.935 0.868 0.917 0.870
CT 0.995 0.992 0.984 0.987 0.990 0.955 0.968 0.991 0.988
EC 0.338 0.312 0.287 0.327 - 0.309 0.299 0.293 0.298
EP 0.988 0.952 0.902 0.993 0.901 0.961 0.942 0.961 0.971
FM 0.680 0.650 0.610 0.638 0.580 0.607 0.580 0.597 0.617
JV 0.990 0.986 0.981 0.987 0.986 0.983 0.974 0.977 0.978
PEMS 0.830 0.173 0.728 0.795 0.778 0.761 0.669 0.804 0.792
SRS 0.596 0.572 0.542 0.564 0.541 0.546 0.507 0.574 0.565
DDG 0.660 0.240 0.400 0.620 0.340 0.280 0.200 0.213 0.624
3 Experiments
Since this work attempts to propose a generic network architecture for time series analysis, we extensively perform experiments over two categories of mainstream time series mining tasks: forecasting and classification.
3.1 Experimental Setup
Table 3: Experimental results w.r.t. studying the effectiveness of deformable patch embedding.
Pointwise Uniform Patchwise DePatch-Conv DePatch-Conv-Conv DePatch-MLP
Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
AWR 0.961 0.961 0.979 0.978 0.981 0.981 0.987 0.987 0.979 0.978
CR 0.958 0.957 0.982 0.981 0.965 0.966 0.986 0.986 0.972 0.972
CT 0.968 0.965 0.995 0.995 0.996 0.996 0.995 0.995 0.994 0.994
EC 0.311 0.238 0.328 0.284 0.356 0.287 0.338 0.249 0.340 0.253
EP 0.986 0.985 0.978 0.978 0.990 0.991 0.988 0.988 0.983 0.983
FM 0.550 0.547 0.663 0.662 0.677 0.676 0.680 0.677 0.647 0.645
JV 0.984 0.983 0.987 0.987 0.985 0.984 0.990 0.990 0.992 0.992
PEMS 0.753 0.746 0.813 0.806 0.817 0.809 0.830 0.823 0.821 0.810
SRS 0.554 0.535 0.578 0.563 0.591 0.585 0.596 0.594 0.574 0.571
DDG 0.527 0.472 0.540 0.515 0.560 0.553 0.660 0.652 0.267 0.160
Time Series Forecasting.
For the task of time series forecasting, we extensively conduct experimental evaluation over $9$ well-known public datasets in our experiments, including ETT (consists of $4$ subsets, i.e., ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, Traffic, Weather, and Illness. To demonstrate the effectiveness of our ConvTimeNet in the prediction of time series, the following advanced approaches are selected as competitive baselines, including convolutional based, such as TimesNetWu et al. (2022), models rooted in Multilayer Perceptrons (MLPs), like N-HitsChallu et al. (2023) and DLinearZeng et al. (2023), and those based on Transformers, including iTransformerLiu et al. (2023), CrossformerZhang and Yan (2022). Furthermore, we have scrutinized top-tier models specifically designed for specific tasks, such as PatchTSTNie et al. (2022) and TiDEDas et al. (2023) for time series forecasting.
Time Series Classification.
For the task of time series classification, our experiments are carried out on a selection of $10$ representative multivariate time series datasets from the renowned UEA archive, which is consistent with recent works Liu et al. (2021a); Wu et al. (2022); Cheng et al. (2023a). To demonstrate the effectiveness of our ConvTimeNet in the classification of time series, we choose the following strong competitive baselines, including self-attention based methods: FormerTime Cheng et al. (2023a), TST Zerveas et al. (2021), Convolutional based: TimesNet Wu et al. (2022), MiniRocket Dempster et al. (2021), TCN Bai et al. (2018), MCDCNN Zheng et al. (2014). In addition, MLP is also employed as our baseline. We do not choose well-known classical distance-based Middlehurst et al. (2021) and shapelet-based Ye and Keogh (2009); He et al. (2023) approaches due to the concern of expensive computation complexity.
For all experimental results, we employed equitable parameter settings. For the evaluation of time series forecasting, We utilize Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the primary metrics for model evaluation. As for classification tasks, we employ Accuracy and the F1 Score as the key indicators for assessment. Noted that all experiments were repeated three times to report the average values. Bold numbers represent the best results and the second best are underlined. Noted that the specific descriptions are provided in the supplement material part.
3.2 Main Results Analysis
Table 1 and Table 2 present the experimental outcomes of ConvTimeNet on the tasks of time series forecasting and classification, respectively. We will subsequently delve into a comprehensive analysis of the performance of both tasks in terms of effectiveness.
Time Series Forecasting.
The data presented in Table 1 unequivocally shows that ConvTimeNet delivers state-of-the-art results, outperforming in over 70% of instances across nine distinct datasets for forecasting tasks. Notably, when compared to PatchTST, ConvTimeNet demonstrates a remarkable improvement in performance, achieving an over 8% increase on the ETTh2 dataset and an impressive 11% enhancement on the illness dataset. Furthermore, when considering MSGNetCai et al. (2023), a hierarchical model of similar structure employed on the Traffic dataset with 862 channels, we observe an exceedingly high memory consumption during training. ConvTimeNet, in stark contrast, maintains an optimal equilibrium between achieving commendable experimental results and maintaining efficiency. These outcomes emphatically underscore ConvTimeNet’s robust capabilities in time series forecasting tasks. However, it is important to note that ConvTimeNet does not consistently yield the best results in every predictive scenario. This observation suggests that there is a potential for further refinement and exploration in enhancing the performance of convolutional models.
Time Series Classification.
In the realm of classification tasks, ConvTimeNet surpasses current state-of-the-art methods in over 80% of cases across ten datasets, a fact depicted in Table 2. Compared to traditional convolutional methods such as MCNN and MCDCNN, ConvTimeNet exhibits a significant improvement in classification accuracy. When juxtaposed with another general purpose network, TimesNet, ConvTimeNet shows a remarkable improvement, exceeding 14% across ten datasets. Note that MiniRocket, a traditional approach, outperformed others on the EC and EP datasets. We guess that the superior performance is likely due to its varying kernel weight design, which facilitates the extraction of more refined features.
Figure 2: Performance comparison between our full-convolution block and Transformer encoder block.
Table 4: Experimental results w.r.t. studying the effectiveness of deep hierarchical architecture. This table only shows the experimental results evaluated by Accuracy while the full results can be found in the appendix part.
Models One Stage Two Stages Three Stages
Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
CT 0.992 0.994 0.995 0.995 0.993 0.995 0.995
FM 0.640 0.680 0.653 0.643 0.660 0.630 0.680
JV 0.990 0.990 0.988 0.987 0.991 0.992 0.990
Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
CR 0.982 0.977 0.982 0.977 0.958 0.972 0.986
AWR 0.984 0.982 0.975 0.982 0.980 0.979 0.987
EC 0.347 0.345 0.342 0.338 0.340 0.337 0.338
PEMS 0.803 0.807 0.829 0.847 0.819 0.807 0.830
DDG 0.573 0.560 0.500 0.580 0.513 0.513 0.660
Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
EP 0.988 0.983 0.981 0.980 0.978 0.976 0.988
SRS2 0.585 0.570 0.589 0.580 0.582 0.589 0.596
3.3 Impacts of Deformable Patch Embedding
The deformable patch embedding plays a vital role in the performance of ConvTimeNet, masterfully tokenizing time series data through its adaptive adjustment of the patch size and position. To rigorously test the effectiveness of this deformable patch mechanism, we embark on a series of experiments across ten different datasets in the realm of classification tasks. These experiments are meticulously designed to compare the outcomes of having pointwise input, implementing a uniform patch, and utilizing three distinct variants of learnable predictor networks in deformable patch. The empirical results, as detailed in Table 3, reveal a significant boost in model performance attributed to the patch operation across most datasets. Particularly noteworthy is its impact on the FM dataset,which records electroencephalogram signals under different finger movements, where it catalyzed a remarkable 20% leap in performance. The prowess of the deformable patch is further highlighted in specific cases, such as the DDG dataset, recording audio information of different animals, where employing a deformable patch configured with a two-layer convolutional network led to a substantial 22% enhancement in model performance. Among the various deformable patch configurations tested, the variant with a two-layer convolutional network emerged as the most effective. In addition, three datasets achieve the best results under the single layer convolution setting, which indicates the deformable patch module can be equipped with different predictors in various scenarios.
3.4 Effectiveness of Fully Convolutional Block
To validate the effectiveness of our designed fully convolutional block, we verify the effectiveness of the block on the forecasting task, due to space limitations, we chose three datasets for the experiment. We replace the fully convolutional block with a traditional transformer encoder block and retain the learnable residual, where the building blocks of our model comprise a minimum of six layers. It is noteworthy that under the condition of setting the transform block to six layers, the model dose not exhibit obvious overfitting phenomena, possibly due to the effect of the learnable residual. This allows for a direct comparison of their forecasting capabilities, as depicted in Figure 2. In a majority of cases, the convolutional block demonstrates greater performance than the transformer encoder block. This effectively validates the superiority of the fully convolutional block in time series prediction tasks. However, on the Exchange dataset, about 5000 exchange rate data, the fully convolution block just slightly surpassing transformer encoder block. This might be attributed to the smaller sample size of the Exchange dataset, thereby causing both blocks to receive insufficient training.
3.5 Impacts of Deep Hierarchical Architecture
To gain a comprehensive understanding of how varying hierarchical structures influence model performance, we conduct an extensive series of experiments. These experiments are conducted across ten datasets dedicated to classification tasks, utilizing senven distinct hierarchical configurations. The objective is to determine the impact of hierarchical depth and complexity on model efficacy. The results, as tabulated in Table 4, reveal a consistent trend: models configured with a three-stage hierarchical structure mostly outperform those with simpler one-stage and two-stage structures. This superior performance underscores the value of a more complex hierarchical arrangement in handling classification tasks. Notably, the experiment results demonstrate that an incremental kernel size approach to hierarchy, where each stage employ larger kernel than the previous, tends to yield more favorable outcomes in most scenarios. One plausible explanation for this observation is that as the model delves deeper and the building block kernel becomes larger, it progressively makes full use of multi-scale information.
3.6 Limitation Discussions
While our proposed ConvTimeNet generally achieves superior performance and reduced computational burden, it is essential to acknowledge its limitations in two key areas. Firstly, for certain datasets, we observed that the hierarchical hyperparameters of ConvTimeNet necessitate meticulous tuning specific to the dataset in question. This process is not only time-consuming but also adds to the implementation costs. To address this challenge, future work could explore the application of neural architecture search Tan and Le (2019) to facilitate the automatic tuning of hyperparameters, thereby alleviating this burden. Secondly, our study only validated the effectiveness of ConvTimeNet within the supervised learning paradigm, without extending experiments to the realm of transfer learning. Recent studies Cheng et al. (2023b) have demonstrated that self-supervised pre-training can significantly enhance the performance of deep learning architectures. Consequently, our approach might not fully realize its potential without the benefits of self-supervised pre-training. Future research could consider integrating self-supervised learning strategies with ConvTimeNet to further enhance its applicability and effectiveness.
4 Conclusion
In this study, we delved into the research question of how to reinvigorate the role of convolutional networks in time series analysis modeling. Specifically, we proposed ConvTimeNet, a deep hierarchical fully convolutional network, which can serve as a versatile backbone network for time series analysis. One key finding of ConvTimeNet is that preserving a deep and hierarchical convolutional network, equipped with modern techniques, can yield superior or competitive performance compared to the prevalent Transformer network and pioneering convolutional model. Extensive experiments conducted on the forecasting and classification of time series fully substantiate its effectiveness. Overall, we hope that ConvTimeNet can serve as an alternative model and encourage the research community to rethink the importance of convolution in time series mining tasks.
Appendix A Related Work
Time series temporal and cross-variable dependency modeling are crucial aspects of time series analysis. In recent years, a variety of modeling techniques have been developed, including those rooted in Convolutional Neural Networks (CNN), Multi-Layer Perceptrons (MLP), and Transformer networks. The CNN-based methods, such as those in Zheng et al. [2014]; Cui et al. [2016], employ sliding convolutional kernels along the temporal dimension to capture temporal dependency. However, these methods have not yielded ideal results in modeling long-range dependencies due to the limited receptive field. On the other hand, MLP-based methods, such as those in Zeng et al. [2023]; Challu et al. [2023], utilize the MLP structure to encode temporal dependencies into the MLP layers. Alternatively, some methods, such as those in Das et al. [2023], can integrate covariate information into the network.
The Transformer network, renowned for its ability to capture long-range dependencies and cross-variable interactions, is particularly appealing for time series analysis. Consequently, numerous Transformer-based methods have been developed. For instance, Autoformer Wu et al. [2021] utilizes a self-correlation mechanism to capture temporal dependencies. Crossformer Zhang and Yan [2022] is another Transformer-based method that introduces a customized dimension segmentation embedding scheme and an explicit cross-variable attention module designed for forecasting tasks. FormerTime Cheng et al. [2023a] and GPTH Liu et al. [2024b], on the other hand, employ a hierarchical structure to capture different-scale temporal dependencies and variable dependencies. Meanwhile, the temporal patching operation Nie et al. [2022] has significantly improved the performance of the model. Despite this, the application of convolutional networks in time series analysis has gradually become less common.
As modern convolutional techniques continue to evolve, a growing number of convolutional methods are being revisited within the community. For instance, the method Ding et al. [2022] propose RepLKNet, which employs a mechanism of re-parameter to circumvent issues associated with expanding the convolutional kernel, thereby addressing the issue of receptive field limitation. Additionally, the TimesNet, introduced recently in Wu et al. [2022], considers two-dimensional temporal variations generated by periodicities, specifically designed for general tasks in time series analysis.
Appendix B Datasets
Table 5: The dataset descriptions are comprehensive. The variable ‘Dim’ represents the number of variables in each dataset. The term ‘Size’ refers to the total number of time steps in the (Train, Validation, Test) set. The term ‘Series Length’ in Time Series Forecasting indicates the future time steps to be predicted, and input series length in Time Series Classification. Each dataset in forecasting task comprises four prediction settings. The term ‘Frequency’ denotes the sampling interval of time steps.
Task Dataset Short Name Dim Series Length Dataset Size Information Frequency
Time Series Forecasting ETTm1, ETTm2 ETTm1, ETTm2 7 (96, 192, 336, 720) (34465, 11521, 11521) Electricity 15min
ETTh1, ETTh2 ETTh1, ETTh2 7 (96, 192, 336, 720) (8545, 2881, 2881) Electricity Hourly
Electricity Electricity 321 (96, 192, 336, 720) (18317, 2633, 5261) Electricity Hourly
Traffic Traffic 862 (96, 192, 336, 720) (12185, 1757, 3509) Transportation Hourly
Weather Weather 21 (96, 192, 336, 720) (36792, 5271, 10540) Weather 10min
Exchange Exchange 8 (96, 192, 336, 720) (5120, 665, 1422) Economy Daily
Illness Ill 7 (24, 36, 48, 60) (617, 74, 170) Medicine Weekly
Time Series Classification ArticularyWordRecognition AWR 9 144 (275,0,300) Motion 200HZ
CharacterTrajectories CT 3 182 (1442,0,1436) Motion 200HZ
Cricket CR 6 1197 (108,0,72) HAR 184HZ
DuckDuckGeese DDG 1345 270 (50,0,50) Audio 44100HZ
Epilepsy EP 3 206 (137,0,138) HAR 16HZ
EthanolConcentration EC 3 1751 (261,0,263) Spectro 0.5nm (of spectrum)
FingerMovements FM 28 50 (316,0,100) EEG 1000HZ
JapaneseVowels JV 12 29 (270,0,370) Audio None
PEMS-SF PEMS 963 144 (267,0,173) Transportation 10 minus
SelfRegulationSCP2 SRS 7 1152 (200,0,180) EEG 256HZ
Figure 3: The pipeline of time series forecasting task.
Figure 4: The pipeline of time series classification task.
Table 6: The key parameters of the experiment, where ‘DW-Kernel Size’ represents the size of the deepwise convolutional kernel in different fully convolutional blocks. ‘Patch Size’ represents the size of the patch window, and ‘Stride’ represents the stride of the patch window.
Task Dataset DW-Kernel Size Patch Size Stride Hidden Size Learning Rate Dropout
Classification ArticularyWordRecognition 19,19,29,29,37,37 32 0.5$\times$×Patch Size 64 0.001 0.1
CharacterTrajectories 7,7,13,13,19,19 32
Cricket 19,19,29,29,37,37 8
DuckDuckGeese 19,19,29,29,37,37 8
Epilepsy 37,37,43,43,53,53 4
EthanolConcentration 19,19,29,29,37,37 32
FingerMovements 7,7,13,13,19,19 64
JapaneseVowels 7,7,13,13,19,19 4
PEMS-SF 19,19,29,29,37,37 16
SelfRegulationSCP2 37,37,43,43,53,53 32
Forecasting ETTh1 9,11,15,21,29,39 32 0.5$\times$×Patch Size 64 [0.0001,0.005] [0.1,0.5]
ETTh2 5,7,9,11,13,15 32
ETTm1 9,11,15,21,29,39 32
ETTm2 11,15,19,25,33,43,55 32
Traffic 15,17,21,27,35,55 32
Weather 9,11,15,19,25,33,43 32
Electricity 11,15,21,29,39,51 32
Exchange Rate 9,11,15,21,29,39 32
Illness 9,11,15,21,29,39 32
In regard to forecasting task, we adhere to the same data processing and train-validation-test set split protocol utilized in TimesNet Wu et al. [2022]. This protocol ensures the strict division of the train, validation, and test datasets based on chronological order, thereby mitigating any data leakage issues. In the time series classification task, we employ the same preprocessing approach as FormerTimeCheng et al. [2023a]. And we select ten datasets, encompassing various tasks such as gesture, action, and audio recognition, medical signal discrimination and others. Detailed information about these datasets is provided in the Table 5.
Time Series Forescasting.
While for time series forecasting, we conduct experiments using nine different datasets, each of which is briefly described as follows: (1) ETT [^2] https://github.com/zhouhaoyi/ETDataset encompasses records of oil temperature and load metrics of electricity transformers. The data spans from July 2016 to July 2018, divided into four sub-datasets. (2) Electricity [^3] https://archive.ics.uci.edu/ml/datasets/ comprises electricity usage data from 321 clients. This data covers the period from July 2016 to July 2019. (3) Exchange [^4] https://github.com/laiguokun/multivariate-time-series-data contains daily exchange rates of eight different nations. The data spans from 1990 to 2016. (4) Traffic [^5] http://pems.dot.ca.gov provides hourly traffic volume data on San Francisco freeways, recorded by 862 sensors from 2015 to 2016. (5) Weather [^6] https://www.bgc-jena.mpg.de/wetter/ consists of 21 weather indicators, including air temperature and humidity. The data was collected at 10-minutes intervals throughout the year 2021. (6) Illness  [^7] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html logs the weekly ratio of patients with influenza-like symptoms against the total patients, collected by the Centers for Disease Control and Prevention of the United States from 2002 to 2021.
In addition, we maintain a fixed length of the lookback window at 336 for ETT (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, Exchange, Traffic and 104 for Illness dataset. The predicted length, varies within the range of 96, 192, 336, 720 and 24, 36, 48, 60 for Illness.
Time Series Classification.
The UEA Bagnall et al. [2018] archive is recognized as one of the most comprehensive benchmarks for multivariate time series analysis. We carefully pick ten datasets that exhibit a variety of characteristics, including differing numbers and lengths of time series samples, as well as class variety. Each of the ten datasets is briefly described as follows: (1) ArticularyWordRecognition contains data collected from multiple native English speakers producing 25 words, to measure the movement of the tongue and lips. (2) CharacterTrajectories captures 2,858 character samples with three dimensions, to identify one of the 20 characters. (3) Cricket records 4 umpires motions while performing 12 cricket signals with 10 repetitions each, aiming to classify the signals based on the accelerometer data. (4) DuckDuckGeese is collected from recordings on www.xenocanto.com, to distinguish different bird species. (5) Epilepsy records data from 6 healthy participants while performing 4 different activities, and the target is to classify those activities. (6) EthanolConcentration records raw spectra of water-and-ethanol solutions in 44 whisky bottles, to determine the alcohol concentration. (7) FingerMovements records a normal subject during a no-feedback session, aiming to classify upcoming movements based on Electroencephalography (EEG) recordings from 28 channels. (8) JapaneseVowels records 9 Japanese-male speakers saying the vowels ‘a’ and ‘e’, to predict the speaker based on the transformed utterances. (9) PEMS-SF describes the occupancy rate of car lanes in the San Francisco Bay area freeways, to classify each observed day into the correct day of the week. (10) SelfRegulationSCP2 records the slow cortical potentials of an Amyptrophic Lateral Sclerosis (ALS) patient, and the task is to predict the direction of cursor movement based on the EEG data.
Appendix C Model Configuration
In both tasks, the Deformable patch slicing module employs two convolution layers as its light weight predictor. The design of the fully convolutional block remains uniform, utilizing one layer of deepwise convolution and two layers of pointwise convolution. Additionally, it incorporates learnable residual and reparameter mechanism. Generally, the small kernel size for reparameter is set to either 3 or 5.
Network for Time Series Forecasting.
For time series forecasting tasks, the pipeline of forward is shown in Figure 3. The input consists of a time series with $C$ variables, each having a specific sequence length $T$. We also employ the series stationarization technique from RevIN Kim et al. [2021] to mitigate the impact of distribution shifts. To ensure that the model can effectively capture multi-scale information and global receptive field, the number of fully convolution blocks is set to six or seven. For the output prediction layer, we use Flatten and Linear layers as output projection layer, aligning with the setup of PatchTST. Denote $X_{out}\in\mathbb{R}^{C\times T_{pl}}$ as output result, where $T_{pl}$ represents the predicted length.
Network for Time Series Classification.
The pipeline of time series classification task is shown in Figure 4. The input data is a time series comprising $C$ variables, each with a distinct sequence length $T$. The number of fully convolution blocks is also generally set to six. For the output layer, we employ max pooling, maintaining consistency with the FormerTime approach Cheng et al. [2023a]. The projection layer, activated with log-softmax, is designed to transform the final representation into the final probability of each class, $X_{out}\in\mathbb{R}^{C_{nc}\times 1}$, where $C_{nc}$ represents the number of class.
Appendix D Implement Details
All experiments are conducted using PyTorch Paszke et al. [2019] on one single NVIDIA 4090 24GB GPU. The model optimization utilizes ADAM Kingma and Ba [2014]. For forecasting task, The number of training epochs is fixed at 10. And we use early stopping to prevent model overfitting and set patience into 3. All other hyperparameters and initialization strategies are either derived from the authors of the original works. If no open parameter script is provided, we will set up the experimental parameters according to the default parameter settings, with a learning rate of 0.0001, batch size of 32, and dropout of 0.2. For classification task, the model optimization utilizes ADAM as well. The number of training epochs is fixed at 200. All the baselines reproduced in this study are implemented based on the configurations outlined in the original paper or the official code.
Hyper-parameter.
To ensure the reproducibility of the experiment, we have listed the hyperparameter settings for two tasks in the Table 6. The main parameters include the number of fully convolutional blocks, the varying kernel size within different blocks, the setting of the hidden size, and the range of adjustment for dropout and learning rate. In addition, we have provided experimental scripts in the open-source code, allowing users to run it directly.
In the classification task, we adopted a three-stages hierarchical structure. Each stage contains two fully convolutional blocks with equal kernel sizes. We use a six-stages or seven-stages hierarchy in forecasting tasks, with each stage containing one fully convolutional block. The trend of kernel changes in the both task is clearly from small to large. Meanwhile, we use different sizes of patch windows for different data sets, and uniformly set the sliding step of the window to half of the size of the patch, ensuring that there is an overlap of half of a continuous patch block. We also verify the effectiveness of this patch setting through experiments.
Appendix E More Experimental Results
E.1 The Ablation of Learnable Residual
Table 7: Experimental results w.r.t. studying the effectiveness of learnable redsidual.
W/ Learnable Redsidual W/O Learnable Residual
Metric Accuracy F1 Score Accuracy F1 Score
AWR 0.987 0.987 0.986 0.986
CR 0.986 0.986 0.949 0.948
CT 0.995 0.995 0.995 0.995
EC 0.338 0.249 0.331 0.266
EP 0.988 0.988 0.978 0.979
FM 0.680 0.677 0.627 0.626
JV 0.990 0.990 0.989 0.990
PEMS 0.830 0.823 0.873 0.870
SRS 0.596 0.594 0.624 0.621
DDG 0.660 0.652 0.467 0.415
To mitigate the risk of overfitting potentially associated with the extensive stacking of fully convolutional blocks, we introduced the concept of learnable residual within the architecture of ConvTimeNet. This feature is specifically utilized to balance the model’s complexity with its learning capacity. To empirically evaluate the impact of learnable residual, we conduct a series of ablation studies across ten distinct datasets, focusing on classification tasks.
The results of these studies are illuminating in Table 7. On average, the incorporation of learnable residual into ConvTimeNet led to a 4% enhancement in model performance. This improvement can likely be attributed to the learnable residual, which can dynamically adjust the contribution of each layer to the final output, enabling the model to effectively capture and represent more complex patterns in the data without succumbing to overfitting.
E.2 Performance Comparision of Varying Layer Depth
Table 8: Experimental results w.r.t. studying the hyperparameter sensitivity of the number of fully convolutional blocks.
Depth 3 Layers 6 Layers 9 Layers
Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
AWR 0.978 0.978 0.987 0.987 0.988 0.988
CR 0.977 0.977 0.986 0.986 0.968 0.967
CT 0.992 0.992 0.995 0.995 0.994 0.994
EC 0.331 0.285 0.338 0.249 0.345 0.282
EP 0.978 0.978 0.988 0.988 0.986 0.986
FM 0.673 0.671 0.680 0.677 0.647 0.647
JV 0.994 0.994 0.990 0.990 0.991 0.991
PEMS 0.825 0.817 0.830 0.823 0.804 0.794
SRS 0.576 0.567 0.596 0.594 0.563 0.562
DDG 0.560 0.523 0.660 0.652 0.630 0.623
Table 9: Experimental results w.r.t. studying the hyperparameter sensitivity of patch size. $P$ denotes the patch size, and $S$ represents the stride of the patch window.
Window Size P=32,S=16 P=32,S=32 P=16,S=8 P=16,S=16 P=8,S=4 P=8,S=8
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.369 0.395 0.375 0.400 0.368 0.393 0.369 0.395 0.369 0.394 0.368 0.393
192 0.406 0.414 0.411 0.418 0.406 0.413 0.406 0.414 0.407 0.414 0.406 0.413
336 0.405 0.420 0.407 0.425 0.405 0.420 0.405 0.422 0.410 0.425 0.405 0.421
720 0.442 0.457 0.441 0.458 0.446 0.459 0.443 0.457 0.450 0.462 0.445 0.458
ETTm1 96 0.292 0.345 0.292 0.345 0.297 0.349 0.297 0.349 0.301 0.352 0.296 0.349
192 0.329 0.368 0.330 0.369 0.336 0.375 0.336 0.375 0.343 0.380 0.333 0.372
336 0.363 0.390 0.365 0.393 0.372 0.398 0.372 0.398 0.376 0.403 0.370 0.399
720 0.427 0.428 0.427 0.427 0.435 0.434 0.435 0.434 0.440 0.437 0.433 0.432
Exchange 96 0.086 0.204 0.086 0.204 0.088 0.207 0.088 0.207 0.088 0.208 0.088 0.207
192 0.184 0.303 0.181 0.301 0.181 0.302 0.181 0.302 0.189 0.307 0.185 0.306
336 0.341 0.421 0.335 0.418 0.343 0.425 0.343 0.425 0.335 0.417 0.339 0.421
720 0.879 0.701 1.014 0.765 0.881 0.704 0.881 0.704 0.890 0.707 0.895 0.709
Table 10: The ablation results of re-parameter mechanism.
W/ Reparameter W/O Reparameter
Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
Setting [7,7,7,7,7,7] [7,7,7,13,13,13] [7,7,13,13,19,19] [7,7,7,7,7,7] [7,7,7,13,13,13] [7,7,13,13,19,19]
CT 0.992 0.991 0.995 0.995 0.995 0.995 0.992 0.991 0.992 0.992 0.991 0.990
FM 0.640 0.632 0.643 0.641 0.680 0.677 0.623 0.619 0.637 0.635 0.630 0.625
JV 0.990 0.989 0.987 0.987 0.990 0.990 0.981 0.980 0.981 0.981 0.982 0.981
Setting [19,19,19,19,19,19] [19,19,19,29,29,29] [19,19,29,29,37,37] [19,19,19,19,19,19] [19,19,19,29,29,29] [19,19,29,29,37,37]
CR 0.982 0.981 0.977 0.976 0.986 0.986 0.977 0.976 0.986 0.986 0.963 0.960
AWR 0.984 0.984 0.982 0.982 0.987 0.987 0.981 0.981 0.979 0.979 0.978 0.978
EC 0.347 0.267 0.338 0.273 0.338 0.249 0.321 0.286 0.321 0.296 0.295 0.274
PEMS 0.803 0.797 0.847 0.843 0.830 0.823 0.846 0.841 0.852 0.845 0.856 0.852
DDG 0.573 0.540 0.580 0.578 0.660 0.652 0.513 0.500 0.567 0.544 0.540 0.509
Setting [37,37,37,37,37,37] [37,37,37,43,43,43] [37,37,43,43,53,53] [37,37,37,37,37,37] [37,37,37,43,43,43] [37,37,43,43,53,53]
EP 0.988 0.988 0.981 0.981 0.988 0.988 0.961 0.959 0.978 0.977 0.966 0.965
SRS 0.585 0.579 0.580 0.562 0.596 0.594 0.567 0.564 0.569 0.547 0.552 0.516
To validate the impact of the number of building blocks, we conducted experiments on ten datasets for classification tasks, varying the number of fully convolutional blocks. The results show in Table 8 indicate that stacking building blocks mostly enhances model performance. However, in most datasets, clear overfitting occurs when the number of blocks reaches nine. This could be attributed to the relatively smaller data volumes in classification tasks, which may hinder the model’s ability to undergo sufficient training.
E.3 Impact of Patch Size
The patch slicing operation is a crucial aspect of ConvTimeNet, as it determines the richness of semantic information in time series data. Therefore, we conduct experiments to assess the sensitivity of patch size on three datasets for forecasting tasks. The experimental results are presented in the Table 9, where $P$ denotes the patch size, and $S$ represents the stride of the patch window. The findings indicate that a bigger patch length of 32 typically achieves superior performance in most cases. Additionally, setting the stride to half of the slice length can also enhance model performance. This improvement might be attributed to the general overlap between patches, which ensures the continuity of time series semantics. Furthermore, capturing more patches within a certain length of time series data ensures a richer semantic understanding.
E.4 The Ablation of Re-parameter Mechanism
Within the fully convolutional block of our model, a re-parameter mechanism is integrated, which not only aids in gathering detailed feature insights with small kernel size but also greatly enable the model to expand the kernel size to be larger. To empirically evaluate the efficacy of this re-parameter mechanism, a series of ablation studies are undertaken across a variety of classification task datasets. As the result show in Table 10. Implementing the re-parameter mechanism resulted in an approximate 6% enhancement in the model’s accuracy for classification tasks. The experimental results thus strongly underscore the pivotal role that the re-parameter mechanism plays in elevating the overall performance of the model. Additionally, it is observed from the data that simply increasing the kernel size does not necessarily enhance the model’s performance in scenarios where the re-parameter mechanism is not utilized.
Figure 5: Model performance comparison in time series forecasting task.
Figure 6: Critical difference diagram over the mean ranks of ConvTimeNet, baseline methods of classification task.
Figure 7: The visualization of forecasting case, generated by various models under the input-336-predict-336 setting, is presented. The black lines represent the ground truth, while the orange lines represent the predicted values.
Figure 8: The visualization of deformable patch operation case. The black line represents the original sequence, the orange dotted line box indicates the uniform patch operation, while the blue dotted line box represents the deformable patch operation.
Table 11: Full results of time series classification task, in terms of Accuracy and F1 score. Noted that -- indicates that the model cannot run due to the issue of being out of memory.
Metric Dataset ConvTimeNet FormerTime TimesNet MiniRocket TST MLP TCN MCNN MCDCNN
Accuracy AWR 0.987 0.978 0.980 0.972 0.980 0.960 0.884 0.977 0.980
CR 0.986 0.917 0.889 0.981 0.898 0.935 0.868 0.917 0.870
CT 0.995 0.992 0.984 0.987 0.990 0.955 0.968 0.991 0.988
EC 0.338 0.312 0.287 0.327 - 0.309 0.299 0.293 0.298
EP 0.988 0.952 0.902 0.993 0.901 0.961 0.942 0.961 0.971
FM 0.680 0.650 0.610 0.638 0.580 0.607 0.580 0.597 0.617
JV 0.990 0.986 0.981 0.987 0.986 0.983 0.974 0.977 0.978
PEMS 0.830 0.173 0.728 0.795 0.778 0.761 0.669 0.804 0.792
SRS 0.596 0.572 0.542 0.564 0.541 0.546 0.507 0.574 0.565
DDG 0.660 0.240 0.400 0.620 0.340 0.280 0.200 0.213 0.624
F1 Score AWR 0.987 0.978 0.980 0.970 0.980 0.959 0.883 0.977 0.980
CR 0.986 0.918 0.886 0.980 0.890 0.929 0.861 0.906 0.863
CT 0.995 0.991 0.983 0.986 0.989 0.949 0.966 0.990 0.987
EC 0.249 0.261 0.181 0.298 - 0.248 0.221 0.234 0.241
EP 0.988 0.950 0.896 0.990 0.899 0.961 0.940 0.960 0.971
FM 0.677 0.628 0.604 0.620 0.568 0.602 0.572 0.586 0.605
JV 0.990 0.984 0.980 0.988 0.986 0.982 0.975 0.974 0.976
PEMS 0.823 0.042 0.720 0.786 0.771 0.747 0.613 0.799 0.788
SRS 0.594 0.579 0.507 0.515 0.487 0.516 0.371 0.573 0.565
DDG 0.652 0.113 0.290 0.622 0.227 0.148 0.067 0.092 0.600
Table 12: Full results of studying the effectiveness of deep hierarchical architecture in terms of Accuracy and F1 score.
Metric Hierarchy One Stage Two Stages Three Stages
Accuracy Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
CT 0.992 0.994 0.995 0.995 0.993 0.995 0.995
FM 0.640 0.680 0.653 0.643 0.660 0.630 0.680
JV 0.990 0.990 0.988 0.987 0.991 0.992 0.990
Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
CR 0.982 0.977 0.982 0.977 0.958 0.972 0.986
AWR 0.984 0.982 0.975 0.982 0.980 0.979 0.987
EC 0.347 0.345 0.342 0.338 0.340 0.337 0.338
PEMS 0.803 0.807 0.829 0.847 0.819 0.807 0.830
DDG 0.573 0.560 0.500 0.580 0.513 0.513 0.660
Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
EP 0.988 0.983 0.981 0.981 0.978 0.976 0.988
SRS 0.585 0.570 0.589 0.580 0.582 0.589 0.596
F1 Score Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
CT 0.991 0.994 0.995 0.995 0.993 0.994 0.995
FM 0.632 0.676 0.651 0.641 0.657 0.629 0.677
JV 0.989 0.990 0.988 0.987 0.993 0.991 0.990
Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
CR 0.981 0.977 0.981 0.976 0.958 0.972 0.986
AWR 0.984 0.982 0.974 0.982 0.980 0.979 0.987
EC 0.267 0.304 0.298 0.273 0.267 0.265 0.249
PEMS 0.797 0.798 0.819 0.843 0.812 0.801 0.823
DDG 0.540 0.527 0.496 0.578 0.521 0.452 0.652
Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
EP 0.988 0.983 0.981 0.981 0.979 0.975 0.988
SRS 0.579 0.569 0.579 0.562 0.569 0.563 0.594
Appendix F Case Study Analysis
Forecasting Case.
We present visual case for time series forecasting in Figure 7. These visual representations provide a clear comparison among various models. The study randomly selects a sample from the test dataset as input and plots the predicted results in all models. It can be intuitively observed from the chart that the predicted trend of ConvTimeNet is more stable compared to DLinearZeng et al. [2023] and NHitsChallu et al. [2023]. Although MSGNetCai et al. [2023] also shows relative stability, the predictied trend of ConvTimeNet is more in line with ground truth.
Classification Case.
As shown in Figure 8, we present the visual case of depatch operation. We randomly selecte two series as input to demonstrate the results of both uniform patch and deformable patch. As can be seen from the figure, the deformable patch more prominently envelops the series peaks completely. This might be attributed to the fact that the deformable patch is more sensitive to series with prominent semantic features.
Appendix G Full Results
In this section, we supplement the full experimental results of two tasks and include relevant analysis figures.
G.1 Time Series Forecasting
Figure 5 illustrates the results of all methods on the time series forecasting task. The statistics in the figure represent the average performance of all methods across four different predicted length settings on each dataset.
G.2 Time Series Classification
Figure 6 illustrates the critical difference diagram of classification task as presented in Demšar [2006]. The complete experimental result, shown in the Table 11, is a supplement to the results of Table 2 in the main text. We also present the full experimental results of effectiveness over the hierarchical structure in Table 12, which is a supplement to Table 4 in the main text.

### v2

@@
 ConvTimeNet: A Deep Hierarchical Fully Convolutional Model for Multivariate Time Series Analysis
-Abstract
-This paper introduces ConvTimeNet, a novel deep hierarchical fully convolutional network designed to serve as a general-purpose model for time series analysis.
-The key design of this network is twofold, designed to overcome the limitations of traditional convolutional networks.
-Firstly, we propose an adaptive segmentation of time series into sub-series level patches, treating these as fundamental modeling units.
-This setting avoids the sparsity semantics associated with raw point-level time steps.
-Secondly, we design a fully convolutional block by skillfully integrating deepwise and pointwise convolution operations, following the advanced building block style employed in Transformer encoders.
-This backbone network allows for the effective capture of both global sequence and cross-variable dependence, as it not only incorporates the advancements of Transformer architecture but also inherits the inherent properties of convolution.
-Furthermore, multi-scale representations of given time series instances can be learned by controlling the kernel size flexibly.
-Extensive experiments are conducted on both time series forecasting and classification tasks.
-The results consistently outperformed strong baselines in most situations in terms of effectiveness.
-The code is publicly available.
-[^1] https://github.com/Mingyue-Cheng/ConvTimeNet.
-1 Introduction
-Time series is a sequence of data points arranged in chronological order.
-Fully analyzing and mining this type of data Roberts et al.
-(2013); Kim et al.
-(2021), such as forecasting evolving trends or classifying the target labels for a given instance, can yield significant gains in various fields, including economics, finance, and biology.
-Over the past decades, numerous efforts Box et al.
-(2015); Liu et al.
-(2021b) have been dedicated to this area.
-Among these efforts, deep learning-based methods Zhang et al.
-(2020); Foumani et al.
-(2023); Liu et al.
-(2024a) have made significant progress due to their ability to model non-linear relationships and eliminate labor-intensive feature engineering.
-Over a significant period in the past, the convolutional network He et al.
-(2016); Zheng et al.
-(2014); Middlehurst et al.
-(2023) has played a crucial role in time series analysis, largely due to its inherent properties that strike an excellent balance between computational efficiency and representation quality.
-Data back to the past years, many representative works Bagnall et al.
-(2017) of time series analysis typically employ convolutional networks as the backbone.
-For instance, temporal convolutional network (TCNBai et al.
-(2018)) and its variants are widely used in modeling temporal variation dependence for the time series forecasting task.
-Furthermore, a large number of works (such as InceptionTimeIsmail Fawaz et al.
-(2020), MiniRocketDempster et al.
-(2021), and MCNNCui et al.
-(2016)) are also proposed by employing convolutional networks to identify informative patterns from given instances in the classification of time series.
-However, with the advent of Transformer networks Vaswani et al.
-(2017); Dosovitskiy et al.
-(2020), the role of convolutional architecture in time series analysis appears to be diminishing.
-Recently, Transformer-style networks Wen et al.
-(2022) have nearly become the dominant role in learning representations of time series, ranging from forecasting Zhou et al.
-(2021); Liu et al.
-(2023) to classification tasks Liu et al.
-(2021a).
-In reality, the primary motivation for widely employing the Transformer network is its utilization of the self-attention mechanism.
-This mechanism enables a more effective capture of long-term sequence dependencies.
-Despite their effectiveness, using Transformer networks for time series analysis also has some areas of incompatibility.
-Firstly, multi-scale representations, a crucial element in time series analysis, are challenging to learn within the vanilla Transformer network.
-This is because the self-attention mechanism can only model fixed-scale representations of sequence input.
-Secondly, unlike other types of sequence data, both temporal variation Xiao et al.
-(2022) and cross-variable dependence modeling Fang et al.
-(2023) are crucial for time series analysis.
-However, the Transformer network lacks sufficient consideration of modeling interactions across variables, which carry crucial information for accurate predictions of time series.
-Thirdly, it usually produces quadratic computation complexity Zhou et al.
-(2021) while attending very long point-wise sequence inputs.
-This is difficult to avoid because a numerical value at each time step can struggle to reveal informative information, resulting in very long sequence input.
-Through careful analysis, we observe that these aforementioned predicaments can largely be avoided by leveraging the inherent properties of convolutional networks.
-Specifically, firstly, the ability to learn varying scale representations is easily achieved by maintaining a hierarchical convolutional network architecture Cui et al.
-(2016).
-Secondly, it is established that the convolutional filter operator naturally complements the modeling of cross-variable relationships Liu et al.
-(2023).
-When it comes to temporal dependence, the receptive field can be progressively expanded by stacking deep convolutional operations, thereby acquiring the capacity for global sequence dependence.
-Thirdly, convolutional neural networks are highly efficient in computation, primarily due to their weight-sharing mechanism Han et al.
-(2021).
-This efficiency distinguishes them from Transformer networks, which are constrained by the curse of quadratic time complexity, which is directly proportional to the square of the input size.
-Along this line, two questions naturally arise: what is the primary obstacle in applying current convolutional networks to time series analysis?
-If these barriers are overcome, can convolutional networks be revolutionized to achieve superior performance again?
-In response to the aforementioned questions, in this work, we propose a novel convolutional network, dubbed ConvTimeNet, designed to serve as a versatile model for time series analysis.
+Abstract.
+Designing effective models for learning time series representations is foundational for time series analysis.
+Many previous works have explored time series representation modeling approaches and have made progress in this area.
+Despite their effectiveness, they lack adaptive perception of local patterns in temporally dependent basic units and fail to capture the multi-scale dependency among these units.
+Instead of relying on prevalent methods centered around self-attention mechanisms, we propose ConvTimeNet, a hierarchical pure convolutional model designed for time series analysis.
+ConvTimeNet introduces a deformable patch layer that adaptively perceives local patterns of temporally dependent basic units in a data-driven manner.
+Based on the extracted local patterns, hierarchical pure convolutional blocks are designed to capture dependency relationships among the representations of basic units at different scales.
+Moreover, a large kernel mechanism is employed to ensure that convolutional blocks can be deeply stacked, thereby achieving a larger receptive field.
+In this way, local patterns and their multi-scale dependencies can be effectively modeled within a single model.
+Extensive experiments comparing a wide range of different types of models demonstrate that pure convolutional models still exhibit strong viability, effectively addressing the aforementioned two challenges and showing superior performance across multiple tasks.
+The code is available for reproducibility [^1] https://github.com/Mingyue-Cheng/ConvTimeNet.
+Time series classification, Deep convolution network
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+1. Introduction
+Time series, sequences of data points arranged in chronological order, are fundamental to various fields due to the valuable insights they provide through analysis and mining (Roberts et al., 2013; Kim et al., 2021).
+In the realm of computer networks, time series analysis plays a pivotal role, especially amid the rapid expansion of the internet infrastructures.
+The ability to analyze and interpret temporal patterns in network data is essential for numerous applications.
+For instance, monitoring network traffic over time aids in detecting anomalies, predicting congestion, and ensuring network security through intrusion detection systems.
+Moreover, time series analysis enables network administrators to proactively address performance issues by forecasting bandwidth utilization and latency fluctuations, thereby enhancing the efficiency and reliability of network operations.
+Over the past decades, numerous efforts (Box et al., 2015; Liu et al., 2021b) have been dedicated to this area.
+Initially, classical statistical methods ARIMA (Zhang, 2003) were predominant because of their mathematical rigor and interpretability.
+However, these approaches often struggled with capturing nonlinear patterns and scaling to large datasets.
+The advent of deep learning introduced data-driven methods such as Long Short-Term Memory (LSTM) networks (Lou et al., 2022) and Gated Recurrent Units (GRUs) (Petneházi, 2019), which modeled sequential data by maintaining hidden states that capture temporal dependencies.
+Despite advancing the field, these models faced challenges like vanishing gradients and computational inefficiency when handling long sequences.
+To address these limitations, the convolutional network (He et al., 2016; Zheng et al., 2014; Middlehurst et al., 2023) has been adapted for time series analysis, largely due to its inherent properties that strike an excellent balance between computational efficiency and representation quality.
+Data from past years shows that many representative works (Bagnall et al., 2017) of time series analysis typically employ convolutional networks as the backbone.
+For instance, TCN (Bai et al., 2018) and its variants are widely used in modeling temporal variation dependence for the time series forecasting task.
+Furthermore, a large number of works (such as InceptionTime (Ismail Fawaz et al., 2020), MiniRocket (Dempster et al., 2021), and MCNN (Cui et al., 2016)) are also proposed by employing convolutional networks to identify informative patterns from given instances in the classification of time series.
+Recently, Transformer-style networks (Wen et al., 2022) have nearly become the dominant role in learning representations of time series, ranging from forecasting (Zhou et al., 2021; Liu et al., 2023) to classification tasks (Liu et al., 2021a).
+Additionally, novel approaches like Mamba (Gu and Dao, 2023) and adaptations of Graph Neural Networks (GNNs) (Jiang et al., 2023) have emerged, offering alternative perspectives in handling time series data.
+However, with the advent of Transformer networks (Vaswani et al., 2017; Dosovitskiy et al., 2020), the role of convolutional architecture in time series analysis appears to be diminishing.
+Despite the effectiveness of these methods, two fundamental challenges remain unaddressed: local temporal pattern extraction and capturing global sequence dependency.
+For the formal, the current approach is to uniformly segment the time series into patches, treating each patch as a local pattern.
+However, this uniform segmentation may not adaptively capture the most informative local patterns, potentially overlooking critical features.
+Besides, solely focusing on these patterns is insufficient for time series analysis.
+This is because, in time series classification, local patterns serve as the basic units of representation, but without considering the dependencies and relationships among them on a global scale, it is impossible to distinguish between different time series classes.
+It is necessary to analyze the dependencies among local patterns from a global perspective.
+Regarding the latter, although numerous methods have been proposed to model global dependency relationships, they often focus solely on these dependencies while neglecting the effective extraction of local patterns.
+In response to the aforementioned questions, in this work, we propose a novel convolutional network, dubbed ConvTimeNet, designed as a versatile model for time series analysis that simultaneously addresses both local pattern extraction and capturing multi-scale repensentation dependency within a single model.
 One of the most distinctive features of ConvTimeNet is its use of pure convolutional operators to learn the representation of given time series data.
@@
 Notably, we propose to divide the raw time series into sequences of patches in a data-driven manner, which can better preserve the semantics of local regions.
-Secondly, on the basis of the building block in Transformer encoder networks, we design a novel fully convolutional block, where deepwise and pointwise convolutions are organized together simultaneously.
+Secondly, we design a novel fully convolutional block where deepwise and pointwise convolutions are organized together simultaneously.
 Particularly, the proposed ConvTimeNet highlights that very deep and hierarchical network architectures are encouraged and very significant in modeling global receptive fields and learning multi-scale representations of given time series instances.
-To demonstrate the effectiveness and generic capacity of ConvTimeNet, we conduct extensive experiments over a variety of time series analysis tasks including forecasting and classification.
+To demonstrate the effectiveness and generic capacity of ConvTimeNet, we conduct extensive experiments over a variety of time series analysis tasks, including forecasting and classification.
 The experimental results show that ConvTimeNet could achieve superior or competitive performance compared to strong baselines, including both advanced Transformer networks and pioneering convolutional models.
-We hope that the proposed ConvTimeNet can serve as an alternative and useful encoder for time series analysis.
-2 A Recipe of Fully Convolutional Network for Time Series Analysis
-Figure 1: Illustration of the newly proposed ConvTimeNet.
+We hope that the proposed ConvTimeNet can serve as an alternative and useful model for time series analysis.
+Main contributions of this paper can be summarized as follows:
+  • We propose ConvTimeNet, a novel fully hierarchical convolutional model that addresses both local pattern extraction and global sequence dependency in time series analysis.
+  • We design a convolutional building block that can be deeply stacked by utilizing depthwise and pointwise convolutions along with a reparameterization mechanism.
+  • We construct extensive experiments to demonstrate that pure convolutional models can outperform strong baselines on multiple tasks, and numerous additional experiments further confirm the effectiveness of our model design.
+2. Related Work
+Time series local pattern extraction and capturing its dependency modeling are crucial aspects of time series analysis.
+Over the years, various approaches have been proposed to model these dependencies, ranging from traditional statistical methods to modern data-driven techniques.
+Early research predominantly focused on statistical methods such as the ARIMA (Zhang, 2003), which constructs autoregressive models and forecasts using moving averages.
+ARIMA effectively captures linear relationships within stationary time series data by combining autoregressive terms (AR), differencing operations (I), and moving average terms (MA) to model temporal dependencies.
+In recent years, deep learning has ushered in a new era of data-driven methods for time series analysis, including those rooted in Convolutional Neural Networks (CNN), Multi-Layer Perceptrons (MLP), and Transformer networks.
+The CNN-based methods, such as those in (Zheng et al., 2014; Cui et al., 2016), employ sliding convolutional kernels along the temporal dimension to capture temporal dependency.
+However, these methods have not yielded ideal results in modeling long-range dependencies due to the limited receptive field.
+On the other hand, MLP-based methods, such as those in (Zeng et al., 2023; Challu et al., 2023), utilize the MLP structure to encode temporal dependencies into the MLP layers.
+Alternatively, some methods, such as those in (Das et al., 2023), can integrate covariate information into the network.
+The Transformer network, renowned for its ability to capture long-range dependencies and cross-variable interactions, is particularly appealing for time series analysis.
+Consequently, numerous Transformer-based methods have been developed.
+For instance, Autoformer (Wu et al., 2021) utilizes a self-correlation mechanism to capture temporal dependencies.
+Crossformer (Zhang and Yan, 2022) is another Transformer-based method that introduces a customized dimension segmentation embedding scheme and an explicit cross-variable attention module designed for forecasting tasks.
+FormerTime (Cheng et al., 2023a) and GPTH (Liu et al., 2024), on the other hand, employ a hierarchical structure to capture different-scale temporal dependencies and variable dependencies.
+Meanwhile, the temporal patching operation (Nie et al., 2022) has significantly improved the performance of the model.
+Despite this, the application of convolutional networks in time series analysis has gradually become less common.
+Recently, the application of Large Language Models (LLMs) to time series analysis (Tan et al., 2024; Chang et al., 2023; Cheng et al., 2024) has attracted significant attention in the research community.
+In the Time-LLM (Jin et al., 2023), the input time series is first tokenized through patching and then aligned with low-dimensional word embeddings using multi-head attention.
+The aligned outputs, combined with embeddings of descriptive statistical features, are then fed into a frozen, pre-trained language model.
+The output representations from the language model are flattened and passed through a linear layer to generate the forecast.
+In this process, the patching employs a uniform segmentation approach.
+As modern convolutional techniques continue to evolve, a growing number of convolutional methods are being revisited within the community.
+For instance, the method (Ding et al., 2022) proposes RepLKNet, which employs a mechanism of re-parameter to circumvent issues associated with expanding the convolutional kernel, thereby addressing the issue of receptive field limitation.
+Additionally, the TimesNet, introduced recently in (Wu et al., 2022), considers two-dimensional temporal variations generated by periodicities, designed for general tasks in time series analysis.
+3. A Recipe of Fully Convolutional Network for Time Series Analysis
+Figure 1. Illustration of the newly proposed ConvTimeNet.
+[…]
+Illustration of the newly proposed ConvTimeNet.
 In this section, we begin by providing a brief overview of the architecture of the newly proposed ConvTimeNet.
@@
 Finally, we demonstrate the strengths of the proposed ConvTimeNet, examining its capabilities from three distinct perspectives.
-2.1 Overall Network Architecture
+3.1. Overall Network Architecture
 As illustrated in Figure 1, the forward pass of ConvTimeNet begins with the input of multivariate time series data.
-Before learning the representation of input data with hidden layers, we develop a deformable patch embedding layer.
-This layer’s primary role is to transform the raw point-wise time series input into a sequence of patch embeddings.
-Specifically, the selection of time series points for each patch is determined in a data-driven manner, facilitated by a predictor layer.
-Subsequently, the transformed time series representation is fed into a stack of full-convolution blocks, with the objective of capturing long-term dependence and learning multi-scale representations.
-Finally, the architecture culminates in a linear layer, designed to perform time series mining tasks such as forecasting or classification of given time series data.
-2.2 Deformable Patch Embedding
+To learn the representations of this time series data, we employ a stack of Fully Convolutional Blocks.
+Leveraging the natural advantages of pure convolution, we vary the sizes of convolutional kernels to deeply model data representations at multiple scales, effectively capturing long-term dependencies.
+Prior to this, and unlike traditional methods that split patches in a fixed manner, we develop a deformable patch embedding layer.
+This data-driven patch splitting approach ensures that patterns within each patch are preserved more reasonably, as the selection of time series points for each patch is determined by a predictor layer.
+Finally, the architecture culminates in a linear layer designed to perform time series mining tasks like forecasting or classification of the given time series data.
+3.2. Adaptively Local Pattern Extraction
 The primary objective of time series analysis is to discern the correlation between data points across different time steps.
@@
 While this approach is effective, we argue that the fixed-size patch splitting may compromise semantics, leading to reduced performance.
-For instance, the same pattern in different time series might have varying sizes, a concept supported by well-known shapelet-based approaches Ye and Keogh (2009).
-Therefore, such strict patch splitting may capture inconsistent information across different instances, potentially leading to sub-optimal results.
-Based on the preceding discussion, we assert that time series analysis has experienced success through patch splitting, yet substantial room for improvement remains.
+For instance, the same pattern in different time series might have varying sizes, a concept supported by well-known shapelet-based approaches (Ye and Keogh, 2009).
+Therefore, strict patch splitting may capture inconsistent information across different instances, likely leading to suboptimal results.
 To address the dilemma mentioned above, we decided to adopt a philosophy of neighboring time points together, selecting specific time points within each patch in a data-driven manner.
-This idea of this approach is also consistent with previous work Dai et al.
-(2017); Chen et al.
-(2021).
+This idea of this approach is also consistent with previous work (Dai et al., 2017; Chen et al., 2021).
 As illustrated in the lower left of Figure 1, we additionally introduced a lightweight predictor to adaptively select time points based on given vanilla features, followed by an embedding projection layer.
@@
 To achieve this goal, the DePatch module comprises three key steps.
-Formally, we denote the input time series as $X\in\mathbb{R}^{C\times T}$, where $C$ and $T$ indicate the number of variables (channels) and sequence lengths respectively.
+Formally, we denote the input time series as $X\in\mathbb{R}^{C\times T}$, where $C$ and $T$ indicate the number of variables (channels) and sequence lengths, respectively.
 Before performing adaptive patch splitting, we first process the raw time series points into $N=\lfloor{\frac{T-P}{S}}\rfloor+2$ fixed-size patches, where the shape of each patch can be denoted by $x\in\mathbb{R}^{P\times C}$.
-We use the padding operation to ensure the uniform split.
+Padding is used to ensure a uniform split.
 For clarity, we omit the superscript ($i$).
@@
 Subsequently, each patch is assigned three key variables: the center position of this patch $x_{c}$, center offset $\delta_{c}$, and scale variation $\delta_{p}$.
-It should be noted that $x_{c}$ is decided by the uniform patch splitting above while the $\delta_{c}$ and $\delta_{p}$ are obtained by borrowing the following computation process:
-$\delta_{c},\delta_{p}=H(g(x)),$ (1)
+It should be noted that $x_{c}$ is decided by the uniform patch splitting above while $\delta_{c}$ and $\delta_{p}$ are obtained by borrowing the following computation process:
+(1) $\delta_{c},\delta_{p}=H(g(x)),$
 where $g(x)$ denotes the feature map obtained from input patch $x$, and $H(\cdot)$ indicates the lightweight predictor function.
@@
 Along this line, the final center of patch $x$ and its scale can be computed by
-$x_{c}^{new}=x_{c}+\delta_{c}$ $,~{}~{}P^{new}=P+2*\delta_{p},$ (2)
-$L=x_{c}^{new}-\frac{P^{new}}{2}$ $,~{}~{}R=x_{c}^{new}+\frac{P^{new}}{2},$
-where $x_{c}^{new}$ denoted as the new center position of the patch, and $P^{new}$ represents the new length of patch.
+(2) $x_{c}^{new}$ $=x_{c}+\delta_{c},P^{new}=P+2\delta_{p},$
+$L$ $=x_{c}^{new}-\frac{P^{new}}{2},R=x_{c}^{new}+\frac{P^{new}}{2}$
+where $x_{c}^{new}$ denoted as the new center position of the patch and $P^{new}$ represents the new length of the patch.
 $L$ and $R$ are denoted as the left and right boundary positions.
-Finally, based on these two positions, we perform sampling on the original sequence with linear interpolation and obtain a new patch, followed by the projection layer transforms the new patch into the embedding space.
+Based on these two positions, we perform sampling on the original sequence with linear interpolation and obtain a new patch, followed by the projection layer transforms the new patch into the embedding space.
 With these efforts, the DePatch module allows for the adjustment of each patch’s position and scale based on the input features adaptively, thereby mitigating the semantic destruction caused by hard splitting.
-2.3 Fully Convolutional Blocks
-To learn high-quality data representation, past methods He et al.
-(2016); Vaswani et al.
-(2017) stacked repeated building blocks, a paradigm that is also very popular in the field of data representation learning.
+3.3. Multi-Scale Representation Dependency Modeling
+To learn high-quality data representation, past methods (He et al., 2016; Vaswani et al., 2017) stacked building blocks, a paradigm that is also very popular in the field of data representation learning.
 Typically, the Transformer block, as a classical design, has been successfully applied in various fields, and these successful applications have inspired methods in the field of time series.
 However, the Transformer block also has some drawbacks.
-Firstly, the traditional Transformer block mainly process fixed-scale input, which is not flexible enough.
-Secondly, the Transformer block is struggling to model the cross-variable dependencies.
-Finally, the computational complexity of the Transformer block is quadratic related to the length of the input sequence.
-Luckily, convolution can effectively address these issues.
-Motivated by this, we propose a fully convolutional block, as shown in the lower right module of Figure 1.
+On one hand, the traditional Transformer block primarily processes fixed-scale input, which limits its flexibility.
+On the other hand, the computational complexity of the Transformer block scales quadratically with the length of the input sequence.
+We propose a fully convolutional block, as shown in the lower right module of Figure 1.
 Specifically, we retain the Transformer block’s innovative architecture, incorporating linear and non-linear transformations.
-The primary distinction lies in our utilization of temporal dependency modeling, achieved through stacking 1D kernel groups, instead of the traditional multi-head attention mechanism.
+The primary distinction lies in our utilization of temporal dependency modeling, achieved through stacking 1D kernel groups instead of the traditional multi-head attention mechanism.
 Practically, we employ deepwise convolution as a 1D kernel group, as it offers efficient computational efficiency.
-Meanwhile we substitute the Feed-Forward Network(FFN) in the Transformer block with a 1x1-sized pointwise convolution to achieve cross-variable dependency modeling.
-The synergistic combination of these two convolutions effectively completing the comprehensive modeling of the temporal and cross variable dependency.
+Meanwhile, we substitute the Feed-Forward Network (FFN) in the Transformer block with a 1x1-sized pointwise convolution to achieve cross-scale dependency modeling.
+The synergistic combination of these two convolutions effectively completed the comprehensive modeling of the temporal and cross-scale dependency.
 To achieve a global receptive field, we increase the depth of the model.
-We discard the traditional residual connection and instead adopt a learnable residual Bachlechner et al.
-(2021), to circumvent the overfitting issue that often arises from an increased model depth.
-We also verifiy the effectiveness of the learnable residual mechanism in our experiments.
-Concurrently, we utilize Batch NormalizationIoffe and Szegedy (2015) in fully convolution block.
+We discard the traditional residual connection and instead adopt a learnable residual (Bachlechner et al., 2021), to circumvent the overfitting issue that often arises from an increased model depth.
+We also verify the effectiveness of the learnable residual mechanism in our experiments.
+Concurrently, we utilize Batch Normalization (Ioffe and Szegedy, 2015) in a fully convolution block.
 Furthermore, we employ larger kernels to expand the receptive field.
 However, in practice, we found that training with large kernels presents a greater challenge.
-This observation is consistent with recent research findingDing et al.
-(2022).
+This observation is consistent with recent research finding (Ding et al., 2022).
 Therefore, we introduce a re-parameterization mechanism to address this issue.
-Due to the flexibility of convolutional networks, we can easily achieve a hierarchical architecture by altering the kernel size.
 We divide the entire deep network architecture into multiple distinct stages, each stage is composed of $K$ fully convolutional blocks.
@@
 We also explored the impact of different configurations of hierarchical structures on the model’s capabilities in the experiment.
-Concretely, denote $F_{\bm{\theta}}^{l}({\cdot})$ as the larger kernel branch and $G_{\bm{\theta}}^{l}({\cdot})$ as the small kernel branch in $l$-th fully convolution block.
-Note that $\theta$ represents the weight of the convolution kernel, it will be updated during gradient backpropagation.
+Concretely, denote $F_{\boldsymbol{\theta}}^{l}(\cdot)$ as the larger kernel branch and $G_{\boldsymbol{\theta}}^{l}(\cdot)$ as the small kernel branch in the $l$-th fully convolutional block.
+Note that $\boldsymbol{\theta}$ represents the weights of the convolution kernels; they will be updated during gradient backpropagation.
 The re-parameterization mechanism is defined as follows:
-${Z}_{DW}^{l}={Gelu}(F_{\bm{\theta}}^{l}({Z^{l-1}})+G_{\bm{\theta}}^{l}({Z^{l-1}})),$ (3)
-where $Z^{l-1}$denote the output of $l-1$ convolutional block, $Z_{DW}^{l}\in\mathcal{\mathbb{R}}^{D\times M}$ represents the output deepwise convolution layer in the $l$-th fully convolutional block.
+(3) ${Z}_{DW}^{l}={Gelu}(F_{\boldsymbol{\theta}}^{l}({Z^{l-1}})+G_{\boldsymbol{\theta}}^{l}({Z^{l-1}})),$
+where $Z^{l-1}$denotes the output of $l-1$ the convolutional block, $Z_{DW}^{l}\in\mathcal{\mathbb{R}}^{D\times M}$ represents the output deepwise convolution layer in the $l$-th fully convolutional block.
 $D$ denoted as the number of hidden dimensions and $M$ represents the new sequence length, which is equal to the number of patches $N$.
-During the inference phase in ConvTimeNet, a sophisticated technique is employed where the weights of the small kernel are zero padded and then combined with those of the larger kernel to form a new, unified convolutional kernel.
-Denote $\alpha$ as the learnable weights for the residual connections in deepwise convolution layer, typically initialized with 00.
-$\hat{Z}_{DW}^{l}=Z^{l-1}+\alpha\times Z_{DW}^{l},$ (4)
+During the inference phase in ConvTimeNet, a sophisticated technique is employed where the weights of the small kernel are zero-padded and then combined with those of the larger kernel to form a unified convolutional kernel.
+Denote $\alpha$ as the learnable weights for the residual connections in the deepwise convolution layer, typically initialized with 00.
+(4) $\hat{Z}_{DW}^{l}=Z^{l-1}+\alpha\times Z_{DW}^{l},$
 where $\hat{Z}_{DW}^{l}$ denote the output of deepwise convolution in $l$-th building block.
@@
 The design of hierarchical and large kernels for ConvTimeNet facilitates the representation learning of varying scales and the creation of a global receptive field.
-2.4 Analysis of the ConvTimeNet
-Through these efforts, a deep and hierarchical convolutional network has been constructed, serving as a general-purpose model for time series analysis.
-It is noteworthy that the proposed ConvTimeNet not only retains the advanced properties of the Transformer encoder but also inherits the pioneering convolutional network.
-Consequently, we explore the relationship between our ConvTimeNet and traditional convolutional networks, as well as Transformer networks.
-ConvTimeNet is equipped with more modern techniques compared to traditional convolutional networks.
-It enhances the semantic information of the input through the use of a deformable patch module and incorporates numerous great designs from the Transformer architecture.
-Furthermore, unlike past convolutional methods in the time series field, we consider the convolution block from operator level and utilize deepwise convolution to enhance the computational efficiency of the model.
-In comparison to the Transformer method, ConvTimeNet foregoes the complex self-attention module and instead constructs a cross-variable dependency model through the collaborative role of deepwise and pointwise convolutions.
-Simultaneously, due to the natural advantages of convolution, a multi-scale architecture can be directly implemented by varying the convolution kernels.
-The hierarchical architecture and large kernel in ConvTimeNet endow it with the capacity to possess a global receptive field.
-Table 1: Experimental results of time series forecasting task evaluated by MSE and MAE.
+3.4. Analysis of the ConvTimeNet
+ConvTimeNet enhances the semantic representation of the input data using the deformable patch module, which adaptively extracts local temporal patterns in a data-driven manner.
+This module directly tackles the issue of local pattern extraction by allowing the model to focus on the most informative segments of the time series rather than relying on uniform segmentation methods that may overlook critical local features.
+By capturing the nuanced variations within the data, ConvTimeNet preserves essential local semantics crucial for accurate representation learning.
+Furthermore, the fully convolutional block in ConvTimeNet is designed with a combination of depthwise and pointwise convolutions, along with a reparameterization mechanism.
+This architecture enables the model to efficiently capture multi-scale pattern representation dependencies.
+The use of large convolutional kernels within a hierarchical framework allows ConvTimeNet to attain a global receptive field without the computational complexity associated with self-attention mechanisms in Transformer networks.
+By unifying the solutions to both local pattern extraction and global sequence dependency capturing within a single model, ConvTimeNet provides a comprehensive framework that addresses these two fundamental challenges in time series analysis.
+This unification is significant because it allows ConvTimeNet to simultaneously extract informative local features through the deformable patch embedding and capture multi-scale representation dependencies using the fully convolutional block.
+By bridging the gap between traditional convolutional networks and Transformer networks, ConvTimeNet achieves superior performance, offering a holistic approach that previous models have not accomplished within a single architecture.
+4. Experiments
+4.1. Experimental Setup
+Time Series Classification.
+For the task of time series classification, our experiments are carried out on a selection of $10$ representative multivariate time series datasets from the renowned UEA archive, which is consistent with recent works (Liu et al., 2021a; Wu et al., 2022; Cheng et al., 2023a).
+To demonstrate the effectiveness of our ConvTimeNet in the classification of time series, we chose the following strong competitive baselines, including self-attention based methods: FormerTime (Cheng et al., 2023a), TST (Zerveas et al., 2021), Convolutional based: TimesNet (Wu et al., 2022), MiniRocket (Dempster et al., 2021), TCN (Bai et al., 2018), MCDCNN (Zheng et al., 2014).
+In addition, MLP is also employed as our baseline.
+We do not choose well-known classical distance-based (Middlehurst et al., 2021) and shapelet-based (Ye and Keogh, 2009; He et al., 2023) approaches due to the concern of computation complexity.
+For the classification task, the model optimization utilizes ADAM as well.
+The number of training epochs is fixed at 200.
+All the baselines reproduced are implemented based on the configurations outlined in the original paper or the official code.
+Time Series Forecasting.
+For the task of time series forecasting, we extensively conduct experimental evaluation over $9$ well-known public datasets in our experiments, including ETT (consists of $4$ subsets, i.e., ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, Traffic, Weather, and Illness.
+To demonstrate the effectiveness of our ConvTimeNet in the prediction of time series, the following advanced approaches are selected as competitive baselines, including convolutional based, such as TimesNet(Wu et al., 2022), models rooted in Multilayer Perceptrons (MLPs), like N-Hits(Challu et al., 2023) and DLinear(Zeng et al., 2023), and those based on Transformers, including iTransformer(Liu et al., 2023), Crossformer(Zhang and Yan, 2022).
+Furthermore, we have scrutinized top-tier models specifically designed for specific tasks, such as PatchTST(Nie et al., 2022) and TiDE(Das et al., 2023) for time series forecasting.
+For all experimental results, we employed equitable parameter settings and they are conducted on one single NVIDIA 4090 24GB GPU.
+For the evaluation of time series forecasting, we utilize Mean Squared Error (MSE) and Mean Absolute Error (MAE) as primary evaluation metrics.
+For classification tasks, we employ Accuracy as the key indicator for assessment.
+Noted that all experiments were repeated three times to report the average values.
+Bold numbers represent the best results, and the second best are underlined.
+Tables in Appendix present the experimental outcomes of ConvTimeNet on the tasks of time series forecasting and classification, respectively.
+4.2. Main Results Analysis
+The performance of ConvTimeNet was evaluated on both time series forecasting and classification tasks, demonstrating its robust capabilities across various datasets.
+For the forecasting task, the number of training epochs is fixed at 10.
+And we use early stopping to prevent model overfitting and set patience to 3.
+All other hyperparameters and initialization strategies are either derived from the authors of the original works.
+If no open parameter script is provided, we will set up the experimental parameters according to the default parameter settings, with a learning rate of 0.0001, batch size of 32, and dropout of 0.2.
+For the classification task, the number of training epochs is fixed at 200.
+All the baselines reproduced in this study are implemented based on the configurations outlined in the original paper or the official code.
+Figure 2. Critical difference diagram over the mean ranks of ConvTimeNet, baseline methods of classification task.
+Figure 3. Model performance comparison in time series forecasting task.
+Time Series Classification.
+In classification tasks, as depicted in Figure 2 and detailed in Appendix B , ConvTimeNet again outperforms current state-of-the-art methods.
+Figure 2 illustrates a critical difference diagram comparing the mean rankings of ConvTimeNet against various baseline methods, with rankings ranging from 1 (best) to 8 (worst).
+ConvTimeNet achieves the top rank, outperforming other models.
+The results in Appendix B show that ConvTimeNet surpasses other methods in over 80% of cases across ten datasets.
+Compared to traditional convolutional methods like MCNN and MCDCNN, ConvTimeNet exhibits significant improvements in classification accuracy.
+When juxtaposed with another general-purpose network, TimesNet, ConvTimeNet shows remarkable improvements, exceeding 14% across the ten datasets.
+It is worth noting that MiniRocket outperformed others on the EC and EP datasets, likely due to its varying kernel weight design, which facilitates the extraction of more refined features.
+Table 1. Experimental results w.r.t. studying the effectiveness of deformable patch embedding.
+Pointwise Uniform Patchwise DePatch-Conv DePatch-Conv-Conv DePatch-MLP
+Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
+AWR 0.961 0.961 0.979 0.978 0.981 0.981 0.987 0.987 0.979 0.978
+CR 0.958 0.957 0.982 0.981 0.965 0.966 0.986 0.986 0.972 0.972
+CT 0.968 0.965 0.995 0.995 0.996 0.996 0.995 0.995 0.994 0.994
+EC 0.311 0.238 0.328 0.284 0.356 0.287 0.338 0.249 0.340 0.253
+EP 0.986 0.985 0.978 0.978 0.990 0.991 0.988 0.988 0.983 0.983
+FM 0.550 0.547 0.663 0.662 0.677 0.676 0.680 0.677 0.647 0.645
+JV 0.984 0.983 0.987 0.987 0.985 0.984 0.990 0.990 0.992 0.992
+PEMS 0.753 0.746 0.813 0.806 0.817 0.809 0.830 0.823 0.821 0.810
+SRS 0.554 0.535 0.578 0.563 0.591 0.585 0.596 0.594 0.574 0.571
+DDG 0.527 0.472 0.540 0.515 0.560 0.553 0.660 0.652 0.267 0.160
+Time Series Forecasting.
+Figure 3 and Appendix A present a comprehensive comparison of ConvTimeNet with other models on forecasting tasks.
+The radar charts in Figure 3 display model performance using MSE on the left and MAE on the right, with different colored lines representing results across multiple datasets such as Traffic, ETT, and Exchange.
+ConvTimeNet consistently demonstrates superior performance on most datasets, particularly achieving significantly lower errors on the ILI and Weather datasets.
+The data in Appendix A further underscores these findings.
+ConvTimeNet delivers state-of-the-art results, outperforming in over 70% of instances across nine distinct datasets.
+Notably, compared to PatchTST, ConvTimeNet shows remarkable improvements, achieving over an 8% increase on the ETTh2 dataset and an impressive 11% enhancement on the illness dataset.
+Additionally, while models like MSGNet (Cai et al., 2023)—a hierarchical model with a similar structure—exhibit exceedingly high memory consumption during training on datasets like Traffic with 862 channels, ConvTimeNet maintains an optimal balance between performance and efficiency.
+These outcomes highlight ConvTimeNet’s robust capabilities in time series forecasting tasks.
+However, it is important to note that ConvTimeNet does not consistently yield the best results in every predictive scenario, suggesting potential for further refinement and exploration to enhance the performance of convolutional models.
+Figure 4. Performance comparison between our full-convolution block and Transformer encoder block.
+4.3. Impacts of Deformable Patch Embedding
+The deformable patch embedding plays a vital role in the performance of ConvTimeNet, masterfully tokenizing time series data through its adaptive adjustment of the patch size and position.
+These adjusted patches serve as local patterns that capture the essential features of the time series.
+To rigorously test the effectiveness of this deformable patch mechanism, we embark on a series of experiments across ten different datasets in the realm of classification tasks.
+These experiments are meticulously designed to compare the outcomes of having pointwise input, implementing a uniform patch, and utilizing three distinct variants of learnable predictor networks in deformable patches.
+The empirical results, as detailed in Table 1, reveal a significant boost in model performance attributed to the patch operation across most datasets.
+Particularly noteworthy is its impact on the FM dataset, which records electroencephalogram signals under different finger movements, where it catalyzed a remarkable 20% leap in performance.
+The prowess of the deformable patch is further highlighted in specific cases, such as the DDG dataset, recording audio information of different animals, where employing a deformable patch configured with a two-layer convolutional network led to a substantial 22% enhancement in model performance.
+Among the various deformable patch configurations tested, the variant with a two-layer convolutional network emerged as the most effective.
+In addition, three datasets achieve the best results under the single layer convolution setting, which indicates the deformable patch can be equipped with different predictors in various scenarios.
+At the same time, the experimental results validate that convolutional neural networks in our model are capable of performing local pattern extraction.
+4.4. Effectiveness of Fully Convolutional Block
+To validate the effectiveness of our designed fully convolutional block, we verify the effectiveness of the block on the forecasting task.
+Due to space limitations, we chose three datasets for the experiment.
+We replace the fully convolutional block with a traditional transformer encoder block and retain the learnable residual, where the building blocks of our model comprise a minimum of six layers.
+It is noteworthy that under the condition of setting the transform block to six layers, the model does not exhibit obvious overfitting phenomena, possibly due to the effect of the learnable residual.
+This allows for a direct comparison of their forecasting capabilities, as depicted in Figure 4.
+In a majority of cases, the fully convolutional block demonstrates greater performance than the transformer encoder block.
+This effectively validates the superiority of the fully convolutional block in time series forecasting tasks.
+However, on the Exchange dataset, with about 5000 exchange rate data points, the fully convolution block just slightly surpasses the transformer encoder block.
+This might be attributed to the smaller sample size of the Exchange dataset, thereby causing both blocks to receive insufficient training.
+4.5. Impacts of Deep Hierarchical Architecture
+Table 2. Experimental results w.r.t. studying the effectiveness of deep hierarchical architecture.
+Models One Stage Two Stages Three Stages
+Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
+CT 0.992 0.994 0.995 0.995 0.993 0.995 0.995
+FM 0.640 0.680 0.653 0.643 0.660 0.630 0.680
+JV 0.990 0.990 0.988 0.987 0.991 0.992 0.990
+Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
+CR 0.982 0.977 0.982 0.977 0.958 0.972 0.986
+AWR 0.984 0.982 0.975 0.982 0.980 0.979 0.987
+EC 0.347 0.345 0.342 0.338 0.340 0.337 0.338
+PEMS 0.803 0.807 0.829 0.847 0.819 0.807 0.830
+DDG 0.573 0.560 0.500 0.580 0.513 0.513 0.660
+Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
+EP 0.988 0.983 0.981 0.980 0.978 0.976 0.988
+SRS2 0.585 0.570 0.589 0.580 0.582 0.589 0.596
+To gain a comprehensive understanding of how varying hierarchical structures influence model performance, we conduct an extensive series of experiments.
+These experiments are conducted across ten datasets dedicated to classification tasks, utilizing seven distinct hierarchical configurations.
+The objective is to determine the impact of hierarchical depth and complexity on model efficacy.
+The results, as tabulated in Table 2, reveal a consistent trend: models configured with a three-stage hierarchical structure mostly outperform those with simpler one-stage and two-stage structures.
+This superior performance underscores the value of a more complex hierarchical arrangement in handling classification tasks.
+Notably, the experiment results demonstrate that an incremental kernel size approach to hierarchy, where each stage employs a larger kernel than the previous, tends to yield more favorable outcomes in most scenarios.
+One plausible explanation for this observation is that as the model delves deeper and the building block kernel becomes larger, it progressively makes full use of multi-scale information.
+This revision emphasizes that with increased depth and larger kernels, the model enhances the richness of interaction between representations at different scales of the temporal basic patterns.
+4.6. The Ablation of Learnable Residual
+To mitigate the risk of overfitting potentially associated with the extensive stacking of fully convolutional blocks, we introduced the concept of learnable residual within the architecture of ConvTimeNet.
+This feature is specifically utilized to balance the model’s complexity with its learning capacity.
+To empirically evaluate the impact of learnable residual, we conduct a series of ablation studies across ten distinct datasets, focusing on classification tasks.
+The results of these studies are illuminating in Table 3.
+On average, the incorporation of learnable residual into ConvTimeNet led to a 4% enhancement in model performance.
+This improvement can likely be attributed to the learnable residual, which can dynamically adjust the contribution of each layer to the final output, enabling the model to effectively capture and represent more complex patterns in the data without succumbing to overfitting.
+Table 3. Experimental results w.r.t. studying the effectiveness of learnable residual.
+W/ Learnable Redsidual W/O Learnable Residual
+Metric Accuracy F1 Score Accuracy F1 Score
+AWR 0.987 0.987 0.986 0.986
+CR 0.986 0.986 0.949 0.948
+CT 0.995 0.995 0.995 0.995
+EC 0.338 0.249 0.331 0.266
+EP 0.988 0.988 0.978 0.979
+FM 0.680 0.677 0.627 0.626
+JV 0.990 0.990 0.989 0.990
+PEMS 0.830 0.823 0.873 0.870
+SRS 0.596 0.594 0.624 0.621
+DDG 0.660 0.652 0.467 0.415
+Figure 5.
+The visualization of the forecasting case, generated by various models under the input-336-predict-336 setting, is presented.
+The black lines represent the ground truth, while the orange lines represent the predicted values.
+4.7. Case Study
+We present visual cases for time series forecasting in Figure 5.
+In each subplot, the black line represents the ground truth values, while the orange line shows the predicted values produced by the corresponding model.
+For this analysis, we chose DLinear, NHiTS, and TiDE as representatives of MLP-based models.
+PatchTST, iTransformer, and CrossFormer as representatives of Transformer-based models and MSGNet and TimesNet as representatives of convolutional models similar to ours.
+These visual representations provide a clear comparison among various models.
+The study randomly selects a sample from the test dataset as input and plots the predicted results in all models.
+It can be intuitively observed from the chart that the predicted trend of ConvTimeNet is more stable compared to DLinear (Zeng et al., 2023) and NHits(Challu et al., 2023).
+Although MSGNet (Cai et al., 2023) also shows relative stability, the predicted trend of ConvTimeNet is more in line with ground truth.
+4.8. Limitation Discussions
+While our proposed ConvTimeNet generally achieves superior performance and reduced computational burden, it is essential to acknowledge its limitations in two key areas.
+Firstly, there is potential for further enhancement in capturing cross-channel dependencies in multivariate time series tasks.
+In scenarios where multiple variables interact over time, comprehending the dependencies and correlations between different channels becomes crucial for accurate modeling and prediction.
+By considering cross-channel interactions starting from the local pattern extraction stage, the model could be refined to more effectively capture the intricate relationships present in multivariate data.
+Designing an extraction module that holistically integrates local features across multiple variables would provide a more comprehensive understanding of the data, potentially leading to improved performance in tasks requiring nuanced inter-variable analysis.
+Secondly, for certain datasets, we observed that the hierarchical hyperparameters of ConvTimeNet necessitate meticulous tuning specific to the dataset in question.
+This process is not only time-consuming but also adds to the implementation costs.
+To address this challenge, future work could explore the application of neural architecture search (Tan and Le, 2019) to facilitate the automatic tuning of hyperparameters, thereby alleviating this burden.
+Thirdly, our study only validated the effectiveness of ConvTimeNet within the supervised learning paradigm without extending experiments to the realm of transfer learning.
+Recent studies (Cheng et al., 2023b) have demonstrated that self-supervised pre-training can significantly enhance the performance of deep learning architectures.
+Consequently, our approach might not fully realize its potential without the benefits of self-supervised pre-training.
+Future research could consider integrating self-supervised learning strategies with ConvTimeNet to further enhance its applicability and effectiveness.
+5. Conclusion
+In this study, we delved into the two major challenges faced in time series data analysis: accurate local pattern extraction and effective global sequence dependency capturing.
+We addressed the research question of how to reinvigorate the role of convolutional networks in time series modeling.
+Specifically, we proposed ConvTimeNet, a deep hierarchical fully convolutional network that serves as a versatile backbone for time series analysis.
+A key finding of our work is that maintaining a deep and hierarchical convolutional architecture, equipped with modern techniques, can yield superior or competitive performance compared to prevalent Transformer networks and pioneering convolutional models.
+Extensive experiments conducted on time series forecasting and classification tasks fully substantiate its effectiveness.
+Overall, we hope that ConvTimeNet can serve as an alternative model and encourage the research community to rethink the importance of convolution in time series mining tasks.
+By effectively unifying local pattern extraction and global dependency modeling within a single framework, ConvTimeNet demonstrates that convolutional networks still hold significant potential for advancing time series analysis.
+Appendix A: Experimental results of time series forecasting task evaluated by MSE and MAE.
 Methods ConvTimeNet PatchTST iTransformer TiDE MSGNet TimesNet Crossformer DLinear NHits
@@
 60 1.741 0.920 2.168 1.009 2.033 1.013 3.981 1.434 3.008 1.157 2.051 0.989 0.390 0.417 2.400 1.116 2.531 1.087
-Table 2: Experimental results of time series classification task evaluated by Accuracy, where full results in terms of both metrics of Accuracy and F1 score is shown in the supplement materials.
-Noted that -- indicates that the model cannot run due to the issue of being out of memory.
+Appendix B: Results of classification task evaluated by Accuracy. -- indicates the model couldn’t run due to insufficient memory.
 Datasets ConvTimeNet FormerTime TimesNet MiniRocket TST MLP TCN MCNN MCDCNN
@@
 DDG 0.660 0.240 0.400 0.620 0.340 0.280 0.200 0.213 0.624
-3 Experiments
-Since this work attempts to propose a generic network architecture for time series analysis, we extensively perform experiments over two categories of mainstream time series mining tasks: forecasting and classification.
-3.1 Experimental Setup
-Table 3: Experimental results w.r.t. studying the effectiveness of deformable patch embedding.
-Pointwise Uniform Patchwise DePatch-Conv DePatch-Conv-Conv DePatch-MLP
-Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
-AWR 0.961 0.961 0.979 0.978 0.981 0.981 0.987 0.987 0.979 0.978
-CR 0.958 0.957 0.982 0.981 0.965 0.966 0.986 0.986 0.972 0.972
-CT 0.968 0.965 0.995 0.995 0.996 0.996 0.995 0.995 0.994 0.994
-EC 0.311 0.238 0.328 0.284 0.356 0.287 0.338 0.249 0.340 0.253
-EP 0.986 0.985 0.978 0.978 0.990 0.991 0.988 0.988 0.983 0.983
-FM 0.550 0.547 0.663 0.662 0.677 0.676 0.680 0.677 0.647 0.645
-JV 0.984 0.983 0.987 0.987 0.985 0.984 0.990 0.990 0.992 0.992
-PEMS 0.753 0.746 0.813 0.806 0.817 0.809 0.830 0.823 0.821 0.810
-SRS 0.554 0.535 0.578 0.563 0.591 0.585 0.596 0.594 0.574 0.571
-DDG 0.527 0.472 0.540 0.515 0.560 0.553 0.660 0.652 0.267 0.160
-Time Series Forecasting.
-For the task of time series forecasting, we extensively conduct experimental evaluation over $9$ well-known public datasets in our experiments, including ETT (consists of $4$ subsets, i.e., ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, Traffic, Weather, and Illness.
-To demonstrate the effectiveness of our ConvTimeNet in the prediction of time series, the following advanced approaches are selected as competitive baselines, including convolutional based, such as TimesNetWu et al.
-(2022), models rooted in Multilayer Perceptrons (MLPs), like N-HitsChallu et al.
-(2023) and DLinearZeng et al.
-(2023), and those based on Transformers, including iTransformerLiu et al.
-(2023), CrossformerZhang and Yan (2022).
-Furthermore, we have scrutinized top-tier models specifically designed for specific tasks, such as PatchTSTNie et al.
-(2022) and TiDEDas et al.
-(2023) for time series forecasting.
-Time Series Classification.
-For the task of time series classification, our experiments are carried out on a selection of $10$ representative multivariate time series datasets from the renowned UEA archive, which is consistent with recent works Liu et al.
-(2021a); Wu et al.
-(2022); Cheng et al.
-(2023a).
-To demonstrate the effectiveness of our ConvTimeNet in the classification of time series, we choose the following strong competitive baselines, including self-attention based methods: FormerTime Cheng et al.
-(2023a), TST Zerveas et al.
-(2021), Convolutional based: TimesNet Wu et al.
-(2022), MiniRocket Dempster et al.
-(2021), TCN Bai et al.
-(2018), MCDCNN Zheng et al.
-(2014).
-In addition, MLP is also employed as our baseline.
-We do not choose well-known classical distance-based Middlehurst et al.
-(2021) and shapelet-based Ye and Keogh (2009); He et al.
-(2023) approaches due to the concern of expensive computation complexity.
-For all experimental results, we employed equitable parameter settings.
-For the evaluation of time series forecasting, We utilize Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the primary metrics for model evaluation.
-As for classification tasks, we employ Accuracy and the F1 Score as the key indicators for assessment.
-Noted that all experiments were repeated three times to report the average values.
-Bold numbers represent the best results and the second best are underlined.
-Noted that the specific descriptions are provided in the supplement material part.
-3.2 Main Results Analysis
-Table 1 and Table 2 present the experimental outcomes of ConvTimeNet on the tasks of time series forecasting and classification, respectively.
-We will subsequently delve into a comprehensive analysis of the performance of both tasks in terms of effectiveness.
-Time Series Forecasting.
-The data presented in Table 1 unequivocally shows that ConvTimeNet delivers state-of-the-art results, outperforming in over 70% of instances across nine distinct datasets for forecasting tasks.
-Notably, when compared to PatchTST, ConvTimeNet demonstrates a remarkable improvement in performance, achieving an over 8% increase on the ETTh2 dataset and an impressive 11% enhancement on the illness dataset.
-Furthermore, when considering MSGNetCai et al.
-(2023), a hierarchical model of similar structure employed on the Traffic dataset with 862 channels, we observe an exceedingly high memory consumption during training.
-ConvTimeNet, in stark contrast, maintains an optimal equilibrium between achieving commendable experimental results and maintaining efficiency.
-These outcomes emphatically underscore ConvTimeNet’s robust capabilities in time series forecasting tasks.
-However, it is important to note that ConvTimeNet does not consistently yield the best results in every predictive scenario.
-This observation suggests that there is a potential for further refinement and exploration in enhancing the performance of convolutional models.
-Time Series Classification.
-In the realm of classification tasks, ConvTimeNet surpasses current state-of-the-art methods in over 80% of cases across ten datasets, a fact depicted in Table 2.
-Compared to traditional convolutional methods such as MCNN and MCDCNN, ConvTimeNet exhibits a significant improvement in classification accuracy.
-When juxtaposed with another general purpose network, TimesNet, ConvTimeNet shows a remarkable improvement, exceeding 14% across ten datasets.
-Note that MiniRocket, a traditional approach, outperformed others on the EC and EP datasets.
-We guess that the superior performance is likely due to its varying kernel weight design, which facilitates the extraction of more refined features.
-Figure 2: Performance comparison between our full-convolution block and Transformer encoder block.
-Table 4: Experimental results w.r.t.
-studying the effectiveness of deep hierarchical architecture.
-This table only shows the experimental results evaluated by Accuracy while the full results can be found in the appendix part.
-Models One Stage Two Stages Three Stages
-Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
-CT 0.992 0.994 0.995 0.995 0.993 0.995 0.995
-FM 0.640 0.680 0.653 0.643 0.660 0.630 0.680
-JV 0.990 0.990 0.988 0.987 0.991 0.992 0.990
-Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
-CR 0.982 0.977 0.982 0.977 0.958 0.972 0.986
-AWR 0.984 0.982 0.975 0.982 0.980 0.979 0.987
-EC 0.347 0.345 0.342 0.338 0.340 0.337 0.338
-PEMS 0.803 0.807 0.829 0.847 0.819 0.807 0.830
-DDG 0.573 0.560 0.500 0.580 0.513 0.513 0.660
-Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
-EP 0.988 0.983 0.981 0.980 0.978 0.976 0.988
-SRS2 0.585 0.570 0.589 0.580 0.582 0.589 0.596
-3.3 Impacts of Deformable Patch Embedding
-The deformable patch embedding plays a vital role in the performance of ConvTimeNet, masterfully tokenizing time series data through its adaptive adjustment of the patch size and position.
-To rigorously test the effectiveness of this deformable patch mechanism, we embark on a series of experiments across ten different datasets in the realm of classification tasks.
-These experiments are meticulously designed to compare the outcomes of having pointwise input, implementing a uniform patch, and utilizing three distinct variants of learnable predictor networks in deformable patch.
-The empirical results, as detailed in Table 3, reveal a significant boost in model performance attributed to the patch operation across most datasets.
-Particularly noteworthy is its impact on the FM dataset,which records electroencephalogram signals under different finger movements, where it catalyzed a remarkable 20% leap in performance.
-The prowess of the deformable patch is further highlighted in specific cases, such as the DDG dataset, recording audio information of different animals, where employing a deformable patch configured with a two-layer convolutional network led to a substantial 22% enhancement in model performance.
-Among the various deformable patch configurations tested, the variant with a two-layer convolutional network emerged as the most effective.
-In addition, three datasets achieve the best results under the single layer convolution setting, which indicates the deformable patch module can be equipped with different predictors in various scenarios.
-3.4 Effectiveness of Fully Convolutional Block
-To validate the effectiveness of our designed fully convolutional block, we verify the effectiveness of the block on the forecasting task, due to space limitations, we chose three datasets for the experiment.
-We replace the fully convolutional block with a traditional transformer encoder block and retain the learnable residual, where the building blocks of our model comprise a minimum of six layers.
-It is noteworthy that under the condition of setting the transform block to six layers, the model dose not exhibit obvious overfitting phenomena, possibly due to the effect of the learnable residual.
-This allows for a direct comparison of their forecasting capabilities, as depicted in Figure 2.
-In a majority of cases, the convolutional block demonstrates greater performance than the transformer encoder block.
-This effectively validates the superiority of the fully convolutional block in time series prediction tasks.
-However, on the Exchange dataset, about 5000 exchange rate data, the fully convolution block just slightly surpassing transformer encoder block.
-This might be attributed to the smaller sample size of the Exchange dataset, thereby causing both blocks to receive insufficient training.
-3.5 Impacts of Deep Hierarchical Architecture
-To gain a comprehensive understanding of how varying hierarchical structures influence model performance, we conduct an extensive series of experiments.
-These experiments are conducted across ten datasets dedicated to classification tasks, utilizing senven distinct hierarchical configurations.
-The objective is to determine the impact of hierarchical depth and complexity on model efficacy.
-The results, as tabulated in Table 4, reveal a consistent trend: models configured with a three-stage hierarchical structure mostly outperform those with simpler one-stage and two-stage structures.
-This superior performance underscores the value of a more complex hierarchical arrangement in handling classification tasks.
-Notably, the experiment results demonstrate that an incremental kernel size approach to hierarchy, where each stage employ larger kernel than the previous, tends to yield more favorable outcomes in most scenarios.
-One plausible explanation for this observation is that as the model delves deeper and the building block kernel becomes larger, it progressively makes full use of multi-scale information.
-3.6 Limitation Discussions
-While our proposed ConvTimeNet generally achieves superior performance and reduced computational burden, it is essential to acknowledge its limitations in two key areas.
-Firstly, for certain datasets, we observed that the hierarchical hyperparameters of ConvTimeNet necessitate meticulous tuning specific to the dataset in question.
-This process is not only time-consuming but also adds to the implementation costs.
-To address this challenge, future work could explore the application of neural architecture search Tan and Le (2019) to facilitate the automatic tuning of hyperparameters, thereby alleviating this burden.
-Secondly, our study only validated the effectiveness of ConvTimeNet within the supervised learning paradigm, without extending experiments to the realm of transfer learning.
-Recent studies Cheng et al.
-(2023b) have demonstrated that self-supervised pre-training can significantly enhance the performance of deep learning architectures.
-Consequently, our approach might not fully realize its potential without the benefits of self-supervised pre-training.
-Future research could consider integrating self-supervised learning strategies with ConvTimeNet to further enhance its applicability and effectiveness.
-4 Conclusion
-In this study, we delved into the research question of how to reinvigorate the role of convolutional networks in time series analysis modeling.
-Specifically, we proposed ConvTimeNet, a deep hierarchical fully convolutional network, which can serve as a versatile backbone network for time series analysis.
-One key finding of ConvTimeNet is that preserving a deep and hierarchical convolutional network, equipped with modern techniques, can yield superior or competitive performance compared to the prevalent Transformer network and pioneering convolutional model.
-Extensive experiments conducted on the forecasting and classification of time series fully substantiate its effectiveness.
-Overall, we hope that ConvTimeNet can serve as an alternative model and encourage the research community to rethink the importance of convolution in time series mining tasks.
-Appendix A Related Work
-Time series temporal and cross-variable dependency modeling are crucial aspects of time series analysis.
-In recent years, a variety of modeling techniques have been developed, including those rooted in Convolutional Neural Networks (CNN), Multi-Layer Perceptrons (MLP), and Transformer networks.
-The CNN-based methods, such as those in Zheng et al.
-[2014]; Cui et al.
-[2016], employ sliding convolutional kernels along the temporal dimension to capture temporal dependency.
-However, these methods have not yielded ideal results in modeling long-range dependencies due to the limited receptive field.
-On the other hand, MLP-based methods, such as those in Zeng et al.
-[2023]; Challu et al.
-[2023], utilize the MLP structure to encode temporal dependencies into the MLP layers.
-Alternatively, some methods, such as those in Das et al.
-[2023], can integrate covariate information into the network.
-The Transformer network, renowned for its ability to capture long-range dependencies and cross-variable interactions, is particularly appealing for time series analysis.
-Consequently, numerous Transformer-based methods have been developed.
-For instance, Autoformer Wu et al.
-[2021] utilizes a self-correlation mechanism to capture temporal dependencies.
-Crossformer Zhang and Yan [2022] is another Transformer-based method that introduces a customized dimension segmentation embedding scheme and an explicit cross-variable attention module designed for forecasting tasks.
-FormerTime Cheng et al.
-[2023a] and GPTH Liu et al.
-[2024b], on the other hand, employ a hierarchical structure to capture different-scale temporal dependencies and variable dependencies.
-Meanwhile, the temporal patching operation Nie et al.
-[2022] has significantly improved the performance of the model.
-Despite this, the application of convolutional networks in time series analysis has gradually become less common.
-As modern convolutional techniques continue to evolve, a growing number of convolutional methods are being revisited within the community.
-For instance, the method Ding et al.
-[2022] propose RepLKNet, which employs a mechanism of re-parameter to circumvent issues associated with expanding the convolutional kernel, thereby addressing the issue of receptive field limitation.
-Additionally, the TimesNet, introduced recently in Wu et al.
-[2022], considers two-dimensional temporal variations generated by periodicities, specifically designed for general tasks in time series analysis.
-Appendix B Datasets
-Table 5: The dataset descriptions are comprehensive.
-The variable ‘Dim’ represents the number of variables in each dataset.
-The term ‘Size’ refers to the total number of time steps in the (Train, Validation, Test) set.
-The term ‘Series Length’ in Time Series Forecasting indicates the future time steps to be predicted, and input series length in Time Series Classification.
-Each dataset in forecasting task comprises four prediction settings.
-The term ‘Frequency’ denotes the sampling interval of time steps.
-Task Dataset Short Name Dim Series Length Dataset Size Information Frequency
-Time Series Forecasting ETTm1, ETTm2 ETTm1, ETTm2 7 (96, 192, 336, 720) (34465, 11521, 11521) Electricity 15min
-ETTh1, ETTh2 ETTh1, ETTh2 7 (96, 192, 336, 720) (8545, 2881, 2881) Electricity Hourly
-Electricity Electricity 321 (96, 192, 336, 720) (18317, 2633, 5261) Electricity Hourly
-Traffic Traffic 862 (96, 192, 336, 720) (12185, 1757, 3509) Transportation Hourly
-Weather Weather 21 (96, 192, 336, 720) (36792, 5271, 10540) Weather 10min
-Exchange Exchange 8 (96, 192, 336, 720) (5120, 665, 1422) Economy Daily
-Illness Ill 7 (24, 36, 48, 60) (617, 74, 170) Medicine Weekly
-Time Series Classification ArticularyWordRecognition AWR 9 144 (275,0,300) Motion 200HZ
-CharacterTrajectories CT 3 182 (1442,0,1436) Motion 200HZ
-Cricket CR 6 1197 (108,0,72) HAR 184HZ
-DuckDuckGeese DDG 1345 270 (50,0,50) Audio 44100HZ
-Epilepsy EP 3 206 (137,0,138) HAR 16HZ
-EthanolConcentration EC 3 1751 (261,0,263) Spectro 0.5nm (of spectrum)
-FingerMovements FM 28 50 (316,0,100) EEG 1000HZ
-JapaneseVowels JV 12 29 (270,0,370) Audio None
-PEMS-SF PEMS 963 144 (267,0,173) Transportation 10 minus
-SelfRegulationSCP2 SRS 7 1152 (200,0,180) EEG 256HZ
-Figure 3: The pipeline of time series forecasting task.
-Figure 4: The pipeline of time series classification task.
-Table 6: The key parameters of the experiment, where ‘DW-Kernel Size’ represents the size of the deepwise convolutional kernel in different fully convolutional blocks.
-‘Patch Size’ represents the size of the patch window, and ‘Stride’ represents the stride of the patch window.
-Task Dataset DW-Kernel Size Patch Size Stride Hidden Size Learning Rate Dropout
-Classification ArticularyWordRecognition 19,19,29,29,37,37 32 0.5$\times$×Patch Size 64 0.001 0.1
-CharacterTrajectories 7,7,13,13,19,19 32
-Cricket 19,19,29,29,37,37 8
-DuckDuckGeese 19,19,29,29,37,37 8
-Epilepsy 37,37,43,43,53,53 4
-EthanolConcentration 19,19,29,29,37,37 32
-FingerMovements 7,7,13,13,19,19 64
-JapaneseVowels 7,7,13,13,19,19 4
-PEMS-SF 19,19,29,29,37,37 16
-SelfRegulationSCP2 37,37,43,43,53,53 32
-Forecasting ETTh1 9,11,15,21,29,39 32 0.5$\times$×Patch Size 64 [0.0001,0.005] [0.1,0.5]
-ETTh2 5,7,9,11,13,15 32
-ETTm1 9,11,15,21,29,39 32
-ETTm2 11,15,19,25,33,43,55 32
-Traffic 15,17,21,27,35,55 32
-Weather 9,11,15,19,25,33,43 32
-Electricity 11,15,21,29,39,51 32
-Exchange Rate 9,11,15,21,29,39 32
-Illness 9,11,15,21,29,39 32
-In regard to forecasting task, we adhere to the same data processing and train-validation-test set split protocol utilized in TimesNet Wu et al.
-[2022].
-This protocol ensures the strict division of the train, validation, and test datasets based on chronological order, thereby mitigating any data leakage issues.
-In the time series classification task, we employ the same preprocessing approach as FormerTimeCheng et al.
-[2023a].
-And we select ten datasets, encompassing various tasks such as gesture, action, and audio recognition, medical signal discrimination and others.
-Detailed information about these datasets is provided in the Table 5.
-Time Series Forescasting.
-While for time series forecasting, we conduct experiments using nine different datasets, each of which is briefly described as follows: (1) ETT [^2] https://github.com/zhouhaoyi/ETDataset encompasses records of oil temperature and load metrics of electricity transformers.
-The data spans from July 2016 to July 2018, divided into four sub-datasets.
-(2) Electricity [^3] https://archive.ics.uci.edu/ml/datasets/ comprises electricity usage data from 321 clients.
-This data covers the period from July 2016 to July 2019.
-(3) Exchange [^4] https://github.com/laiguokun/multivariate-time-series-data contains daily exchange rates of eight different nations.
-The data spans from 1990 to 2016.
-(4) Traffic [^5] http://pems.dot.ca.gov provides hourly traffic volume data on San Francisco freeways, recorded by 862 sensors from 2015 to 2016.
-(5) Weather [^6] https://www.bgc-jena.mpg.de/wetter/ consists of 21 weather indicators, including air temperature and humidity.
-The data was collected at 10-minutes intervals throughout the year 2021.
-(6) Illness  [^7] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html logs the weekly ratio of patients with influenza-like symptoms against the total patients, collected by the Centers for Disease Control and Prevention of the United States from 2002 to 2021.
-In addition, we maintain a fixed length of the lookback window at 336 for ETT (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, Exchange, Traffic and 104 for Illness dataset.
-The predicted length, varies within the range of 96, 192, 336, 720 and 24, 36, 48, 60 for Illness.
-Time Series Classification.
-The UEA Bagnall et al.
-[2018] archive is recognized as one of the most comprehensive benchmarks for multivariate time series analysis.
-We carefully pick ten datasets that exhibit a variety of characteristics, including differing numbers and lengths of time series samples, as well as class variety.
-Each of the ten datasets is briefly described as follows: (1) ArticularyWordRecognition contains data collected from multiple native English speakers producing 25 words, to measure the movement of the tongue and lips.
-(2) CharacterTrajectories captures 2,858 character samples with three dimensions, to identify one of the 20 characters.
-(3) Cricket records 4 umpires motions while performing 12 cricket signals with 10 repetitions each, aiming to classify the signals based on the accelerometer data.
-(4) DuckDuckGeese is collected from recordings on www.xenocanto.com, to distinguish different bird species.
-(5) Epilepsy records data from 6 healthy participants while performing 4 different activities, and the target is to classify those activities.
-(6) EthanolConcentration records raw spectra of water-and-ethanol solutions in 44 whisky bottles, to determine the alcohol concentration.
-(7) FingerMovements records a normal subject during a no-feedback session, aiming to classify upcoming movements based on Electroencephalography (EEG) recordings from 28 channels.
-(8) JapaneseVowels records 9 Japanese-male speakers saying the vowels ‘a’ and ‘e’, to predict the speaker based on the transformed utterances.
-(9) PEMS-SF describes the occupancy rate of car lanes in the San Francisco Bay area freeways, to classify each observed day into the correct day of the week.
-(10) SelfRegulationSCP2 records the slow cortical potentials of an Amyptrophic Lateral Sclerosis (ALS) patient, and the task is to predict the direction of cursor movement based on the EEG data.
-Appendix C Model Configuration
-In both tasks, the Deformable patch slicing module employs two convolution layers as its light weight predictor.
-The design of the fully convolutional block remains uniform, utilizing one layer of deepwise convolution and two layers of pointwise convolution.
-Additionally, it incorporates learnable residual and reparameter mechanism.
-Generally, the small kernel size for reparameter is set to either 3 or 5.
-Network for Time Series Forecasting.
-For time series forecasting tasks, the pipeline of forward is shown in Figure 3.
-The input consists of a time series with $C$ variables, each having a specific sequence length $T$.
-We also employ the series stationarization technique from RevIN Kim et al.
-[2021] to mitigate the impact of distribution shifts.
-To ensure that the model can effectively capture multi-scale information and global receptive field, the number of fully convolution blocks is set to six or seven.
-For the output prediction layer, we use Flatten and Linear layers as output projection layer, aligning with the setup of PatchTST.
-Denote $X_{out}\in\mathbb{R}^{C\times T_{pl}}$ as output result, where $T_{pl}$ represents the predicted length.
-Network for Time Series Classification.
-The pipeline of time series classification task is shown in Figure 4.
-The input data is a time series comprising $C$ variables, each with a distinct sequence length $T$.
-The number of fully convolution blocks is also generally set to six.
-For the output layer, we employ max pooling, maintaining consistency with the FormerTime approach Cheng et al.
-[2023a].
-The projection layer, activated with log-softmax, is designed to transform the final representation into the final probability of each class, $X_{out}\in\mathbb{R}^{C_{nc}\times 1}$, where $C_{nc}$ represents the number of class.
-Appendix D Implement Details
-All experiments are conducted using PyTorch Paszke et al.
-[2019] on one single NVIDIA 4090 24GB GPU.
-The model optimization utilizes ADAM Kingma and Ba [2014].
-For forecasting task, The number of training epochs is fixed at 10.
-And we use early stopping to prevent model overfitting and set patience into 3.
-All other hyperparameters and initialization strategies are either derived from the authors of the original works.
-If no open parameter script is provided, we will set up the experimental parameters according to the default parameter settings, with a learning rate of 0.0001, batch size of 32, and dropout of 0.2.
-For classification task, the model optimization utilizes ADAM as well.
-The number of training epochs is fixed at 200.
-All the baselines reproduced in this study are implemented based on the configurations outlined in the original paper or the official code.
-Hyper-parameter.
-To ensure the reproducibility of the experiment, we have listed the hyperparameter settings for two tasks in the Table 6.
-The main parameters include the number of fully convolutional blocks, the varying kernel size within different blocks, the setting of the hidden size, and the range of adjustment for dropout and learning rate.
-In addition, we have provided experimental scripts in the open-source code, allowing users to run it directly.
-In the classification task, we adopted a three-stages hierarchical structure.
-Each stage contains two fully convolutional blocks with equal kernel sizes.
-We use a six-stages or seven-stages hierarchy in forecasting tasks, with each stage containing one fully convolutional block.
-The trend of kernel changes in the both task is clearly from small to large.
-Meanwhile, we use different sizes of patch windows for different data sets, and uniformly set the sliding step of the window to half of the size of the patch, ensuring that there is an overlap of half of a continuous patch block.
-We also verify the effectiveness of this patch setting through experiments.
-Appendix E More Experimental Results
-E.1 The Ablation of Learnable Residual
-Table 7: Experimental results w.r.t. studying the effectiveness of learnable redsidual.
-W/ Learnable Redsidual W/O Learnable Residual
-Metric Accuracy F1 Score Accuracy F1 Score
-AWR 0.987 0.987 0.986 0.986
-CR 0.986 0.986 0.949 0.948
-CT 0.995 0.995 0.995 0.995
-EC 0.338 0.249 0.331 0.266
-EP 0.988 0.988 0.978 0.979
-FM 0.680 0.677 0.627 0.626
-JV 0.990 0.990 0.989 0.990
-PEMS 0.830 0.823 0.873 0.870
-SRS 0.596 0.594 0.624 0.621
-DDG 0.660 0.652 0.467 0.415
-To mitigate the risk of overfitting potentially associated with the extensive stacking of fully convolutional blocks, we introduced the concept of learnable residual within the architecture of ConvTimeNet.
-This feature is specifically utilized to balance the model’s complexity with its learning capacity.
-To empirically evaluate the impact of learnable residual, we conduct a series of ablation studies across ten distinct datasets, focusing on classification tasks.
-The results of these studies are illuminating in Table 7.
-On average, the incorporation of learnable residual into ConvTimeNet led to a 4% enhancement in model performance.
-This improvement can likely be attributed to the learnable residual, which can dynamically adjust the contribution of each layer to the final output, enabling the model to effectively capture and represent more complex patterns in the data without succumbing to overfitting.
-E.2 Performance Comparision of Varying Layer Depth
-Table 8: Experimental results w.r.t. studying the hyperparameter sensitivity of the number of fully convolutional blocks.
-Depth 3 Layers 6 Layers 9 Layers
-Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
-AWR 0.978 0.978 0.987 0.987 0.988 0.988
-CR 0.977 0.977 0.986 0.986 0.968 0.967
-CT 0.992 0.992 0.995 0.995 0.994 0.994
-EC 0.331 0.285 0.338 0.249 0.345 0.282
-EP 0.978 0.978 0.988 0.988 0.986 0.986
-FM 0.673 0.671 0.680 0.677 0.647 0.647
-JV 0.994 0.994 0.990 0.990 0.991 0.991
-PEMS 0.825 0.817 0.830 0.823 0.804 0.794
-SRS 0.576 0.567 0.596 0.594 0.563 0.562
-DDG 0.560 0.523 0.660 0.652 0.630 0.623
-Table 9: Experimental results w.r.t. studying the hyperparameter sensitivity of patch size. $P$ denotes the patch size, and $S$ represents the stride of the patch window.
-Window Size P=32,S=16 P=32,S=32 P=16,S=8 P=16,S=16 P=8,S=4 P=8,S=8
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 96 0.369 0.395 0.375 0.400 0.368 0.393 0.369 0.395 0.369 0.394 0.368 0.393
-192 0.406 0.414 0.411 0.418 0.406 0.413 0.406 0.414 0.407 0.414 0.406 0.413
-336 0.405 0.420 0.407 0.425 0.405 0.420 0.405 0.422 0.410 0.425 0.405 0.421
-720 0.442 0.457 0.441 0.458 0.446 0.459 0.443 0.457 0.450 0.462 0.445 0.458
-ETTm1 96 0.292 0.345 0.292 0.345 0.297 0.349 0.297 0.349 0.301 0.352 0.296 0.349
-192 0.329 0.368 0.330 0.369 0.336 0.375 0.336 0.375 0.343 0.380 0.333 0.372
-336 0.363 0.390 0.365 0.393 0.372 0.398 0.372 0.398 0.376 0.403 0.370 0.399
-720 0.427 0.428 0.427 0.427 0.435 0.434 0.435 0.434 0.440 0.437 0.433 0.432
-Exchange 96 0.086 0.204 0.086 0.204 0.088 0.207 0.088 0.207 0.088 0.208 0.088 0.207
-192 0.184 0.303 0.181 0.301 0.181 0.302 0.181 0.302 0.189 0.307 0.185 0.306
-336 0.341 0.421 0.335 0.418 0.343 0.425 0.343 0.425 0.335 0.417 0.339 0.421
-720 0.879 0.701 1.014 0.765 0.881 0.704 0.881 0.704 0.890 0.707 0.895 0.709
-Table 10: The ablation results of re-parameter mechanism.
-W/ Reparameter W/O Reparameter
-Metric Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
-Setting [7,7,7,7,7,7] [7,7,7,13,13,13] [7,7,13,13,19,19] [7,7,7,7,7,7] [7,7,7,13,13,13] [7,7,13,13,19,19]
-CT 0.992 0.991 0.995 0.995 0.995 0.995 0.992 0.991 0.992 0.992 0.991 0.990
-FM 0.640 0.632 0.643 0.641 0.680 0.677 0.623 0.619 0.637 0.635 0.630 0.625
-JV 0.990 0.989 0.987 0.987 0.990 0.990 0.981 0.980 0.981 0.981 0.982 0.981
-Setting [19,19,19,19,19,19] [19,19,19,29,29,29] [19,19,29,29,37,37] [19,19,19,19,19,19] [19,19,19,29,29,29] [19,19,29,29,37,37]
-CR 0.982 0.981 0.977 0.976 0.986 0.986 0.977 0.976 0.986 0.986 0.963 0.960
-AWR 0.984 0.984 0.982 0.982 0.987 0.987 0.981 0.981 0.979 0.979 0.978 0.978
-EC 0.347 0.267 0.338 0.273 0.338 0.249 0.321 0.286 0.321 0.296 0.295 0.274
-PEMS 0.803 0.797 0.847 0.843 0.830 0.823 0.846 0.841 0.852 0.845 0.856 0.852
-DDG 0.573 0.540 0.580 0.578 0.660 0.652 0.513 0.500 0.567 0.544 0.540 0.509
-Setting [37,37,37,37,37,37] [37,37,37,43,43,43] [37,37,43,43,53,53] [37,37,37,37,37,37] [37,37,37,43,43,43] [37,37,43,43,53,53]
-EP 0.988 0.988 0.981 0.981 0.988 0.988 0.961 0.959 0.978 0.977 0.966 0.965
-SRS 0.585 0.579 0.580 0.562 0.596 0.594 0.567 0.564 0.569 0.547 0.552 0.516
-To validate the impact of the number of building blocks, we conducted experiments on ten datasets for classification tasks, varying the number of fully convolutional blocks.
-The results show in Table 8 indicate that stacking building blocks mostly enhances model performance.
-However, in most datasets, clear overfitting occurs when the number of blocks reaches nine.
-This could be attributed to the relatively smaller data volumes in classification tasks, which may hinder the model’s ability to undergo sufficient training.
-E.3 Impact of Patch Size
-The patch slicing operation is a crucial aspect of ConvTimeNet, as it determines the richness of semantic information in time series data.
-Therefore, we conduct experiments to assess the sensitivity of patch size on three datasets for forecasting tasks.
-The experimental results are presented in the Table 9, where $P$ denotes the patch size, and $S$ represents the stride of the patch window.
-The findings indicate that a bigger patch length of 32 typically achieves superior performance in most cases.
-Additionally, setting the stride to half of the slice length can also enhance model performance.
-This improvement might be attributed to the general overlap between patches, which ensures the continuity of time series semantics.
-Furthermore, capturing more patches within a certain length of time series data ensures a richer semantic understanding.
-E.4 The Ablation of Re-parameter Mechanism
-Within the fully convolutional block of our model, a re-parameter mechanism is integrated, which not only aids in gathering detailed feature insights with small kernel size but also greatly enable the model to expand the kernel size to be larger.
-To empirically evaluate the efficacy of this re-parameter mechanism, a series of ablation studies are undertaken across a variety of classification task datasets.
-As the result show in Table 10.
-Implementing the re-parameter mechanism resulted in an approximate 6% enhancement in the model’s accuracy for classification tasks.
-The experimental results thus strongly underscore the pivotal role that the re-parameter mechanism plays in elevating the overall performance of the model.
-Additionally, it is observed from the data that simply increasing the kernel size does not necessarily enhance the model’s performance in scenarios where the re-parameter mechanism is not utilized.
-Figure 5: Model performance comparison in time series forecasting task.
-Figure 6: Critical difference diagram over the mean ranks of ConvTimeNet, baseline methods of classification task.
-Figure 7: The visualization of forecasting case, generated by various models under the input-336-predict-336 setting, is presented.
-The black lines represent the ground truth, while the orange lines represent the predicted values.
-Figure 8: The visualization of deformable patch operation case.
-The black line represents the original sequence, the orange dotted line box indicates the uniform patch operation, while the blue dotted line box represents the deformable patch operation.
-Table 11: Full results of time series classification task, in terms of Accuracy and F1 score. Noted that -- indicates that the model cannot run due to the issue of being out of memory.
-Metric Dataset ConvTimeNet FormerTime TimesNet MiniRocket TST MLP TCN MCNN MCDCNN
-Accuracy AWR 0.987 0.978 0.980 0.972 0.980 0.960 0.884 0.977 0.980
-CR 0.986 0.917 0.889 0.981 0.898 0.935 0.868 0.917 0.870
-CT 0.995 0.992 0.984 0.987 0.990 0.955 0.968 0.991 0.988
-EC 0.338 0.312 0.287 0.327 - 0.309 0.299 0.293 0.298
-EP 0.988 0.952 0.902 0.993 0.901 0.961 0.942 0.961 0.971
-FM 0.680 0.650 0.610 0.638 0.580 0.607 0.580 0.597 0.617
-JV 0.990 0.986 0.981 0.987 0.986 0.983 0.974 0.977 0.978
-PEMS 0.830 0.173 0.728 0.795 0.778 0.761 0.669 0.804 0.792
-SRS 0.596 0.572 0.542 0.564 0.541 0.546 0.507 0.574 0.565
-DDG 0.660 0.240 0.400 0.620 0.340 0.280 0.200 0.213 0.624
-F1 Score AWR 0.987 0.978 0.980 0.970 0.980 0.959 0.883 0.977 0.980
-CR 0.986 0.918 0.886 0.980 0.890 0.929 0.861 0.906 0.863
-CT 0.995 0.991 0.983 0.986 0.989 0.949 0.966 0.990 0.987
-EC 0.249 0.261 0.181 0.298 - 0.248 0.221 0.234 0.241
-EP 0.988 0.950 0.896 0.990 0.899 0.961 0.940 0.960 0.971
-FM 0.677 0.628 0.604 0.620 0.568 0.602 0.572 0.586 0.605
-JV 0.990 0.984 0.980 0.988 0.986 0.982 0.975 0.974 0.976
-PEMS 0.823 0.042 0.720 0.786 0.771 0.747 0.613 0.799 0.788
-SRS 0.594 0.579 0.507 0.515 0.487 0.516 0.371 0.573 0.565
-DDG 0.652 0.113 0.290 0.622 0.227 0.148 0.067 0.092 0.600
-Table 12: Full results of studying the effectiveness of deep hierarchical architecture in terms of Accuracy and F1 score.
-Metric Hierarchy One Stage Two Stages Three Stages
-Accuracy Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
-CT 0.992 0.994 0.995 0.995 0.993 0.995 0.995
-FM 0.640 0.680 0.653 0.643 0.660 0.630 0.680
-JV 0.990 0.990 0.988 0.987 0.991 0.992 0.990
-Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
-CR 0.982 0.977 0.982 0.977 0.958 0.972 0.986
-AWR 0.984 0.982 0.975 0.982 0.980 0.979 0.987
-EC 0.347 0.345 0.342 0.338 0.340 0.337 0.338
-PEMS 0.803 0.807 0.829 0.847 0.819 0.807 0.830
-DDG 0.573 0.560 0.500 0.580 0.513 0.513 0.660
-Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
-EP 0.988 0.983 0.981 0.981 0.978 0.976 0.988
-SRS 0.585 0.570 0.589 0.580 0.582 0.589 0.596
-F1 Score Setting [7,7,7,7,7,7] [13,13,13,13,13,13] [19,19,19,19,19,19] [7,7,7,13,13,13] [7,7,7,19,19,19] [13,13,13,19,19,19] [7,7,13,13,19,19]
-CT 0.991 0.994 0.995 0.995 0.993 0.994 0.995
-FM 0.632 0.676 0.651 0.641 0.657 0.629 0.677
-JV 0.989 0.990 0.988 0.987 0.993 0.991 0.990
-Setting [19,19,19,19,19,19] [29,29,29,29,29,29] [37,37,37,37,37,37] [19,19,19,29,29,29] [19,19,19,37,37,37] [29,29,29,37,37,.37] [19,19,29,29,37,37]
-CR 0.981 0.977 0.981 0.976 0.958 0.972 0.986
-AWR 0.984 0.982 0.974 0.982 0.980 0.979 0.987
-EC 0.267 0.304 0.298 0.273 0.267 0.265 0.249
-PEMS 0.797 0.798 0.819 0.843 0.812 0.801 0.823
-DDG 0.540 0.527 0.496 0.578 0.521 0.452 0.652
-Setting [37,37,37,37,37,37] [43,43,43,43,43,43] [53,53,53,53,53,53] [37,37,37,43,43,43] [37,37,37,53,53,53] [43,43,43,53,53,53] [37,37,43,43,53,53]
-EP 0.988 0.983 0.981 0.981 0.979 0.975 0.988
-SRS 0.579 0.569 0.579 0.562 0.569 0.563 0.594
-Appendix F Case Study Analysis
-Forecasting Case.
-We present visual case for time series forecasting in Figure 7.
-These visual representations provide a clear comparison among various models.
-The study randomly selects a sample from the test dataset as input and plots the predicted results in all models.
-It can be intuitively observed from the chart that the predicted trend of ConvTimeNet is more stable compared to DLinearZeng et al.
-[2023] and NHitsChallu et al.
-[2023].
-Although MSGNetCai et al.
-[2023] also shows relative stability, the predictied trend of ConvTimeNet is more in line with ground truth.
-Classification Case.
-As shown in Figure 8, we present the visual case of depatch operation.
-We randomly selecte two series as input to demonstrate the results of both uniform patch and deformable patch.
-As can be seen from the figure, the deformable patch more prominently envelops the series peaks completely.
-This might be attributed to the fact that the deformable patch is more sensitive to series with prominent semantic features.
-Appendix G Full Results
-In this section, we supplement the full experimental results of two tasks and include relevant analysis figures.
-G.1 Time Series Forecasting
-Figure 5 illustrates the results of all methods on the time series forecasting task.
-The statistics in the figure represent the average performance of all methods across four different predicted length settings on each dataset.
-G.2 Time Series Classification
-Figure 6 illustrates the critical difference diagram of classification task as presented in Demšar [2006].
-The complete experimental result, shown in the Table 11, is a supplement to the results of Table 2 in the main text.
-We also present the full experimental results of effectiveness over the hierarchical structure in Table 12, which is a supplement to Table 4 in the main text.